### v2

# Dynamics of the general $Q$-tensor model interacting with a rigid body

Abstract.
In this article, the fluid-rigid body interaction problem of nematic liquid crystals described by the general Beris-Edwards $Q$-tensor model is studied. It is proved first that the total energy of this problem decreases in time. The associated mathematical problem is a quasilinear mixed-order system with moving boundary. After the transformation to a fixed domain, a monolithic approach based on the added mass operator and lifting arguments is employed to establish the maximal $\mathrm{L}^{p}$-regularity of the linearized problem in an anisotropic ground space. This paves the way for the local strong well-posedness for large data and global strong well-posedness for small data of the interaction problem.
Key words and phrases:
Liquid crystal-rigid body interaction, $Q$-tensor model, strong solutions, added mass operator, global well-posedness, energy decay, exponential convergence to equilibria
2020 Mathematics Subject Classification:
35Q35, 76A15, 74F10, 76D03, 35K59
1. Introduction
Nematic liquid crystals are described by the Doi-Onsager, the Ericksen-Leslie and the Landau-de Gennes theory. While Ericksen-Leslie models are vector models, $Q$-tensor models including the Landau-de Gennes theory rely on symmetric, traceless $3\times 3$-matrices $Q$ to model the biaxial alignment of molecules. The evolution of $Q$ is influenced by the free energy of the molecules and aspects such as tumbling and alignment effects. For further background, we refer to the monographs of Virga [45] and Sonnet and Virga [40] or the survey articles of Ball [5], Hieber and Prüss [23], Wang, Zhang and Zhang [46] and Zarnescu [48].
In this work, we investigate the fluid-rigid body interaction problem of the general Beris-Edwards $Q$-tensor model [7]. More precisely, we show the local strong well-posedness for large data and the global strong well-posedness for small data of the interaction problem. In particular, in the underlying $Q$-tensor model, we consider an arbitrary ratio of tumbling and alignment effects $\xi\in\mathbb{R}$. For the fluid velocity $u$ and pressure $\pi$ and the molecular orientation of the liquid crystals $Q$, taking values in the space $\mathbb{S}_{0}^{3}$ of symmetric and traceless $3\times 3$-matrices, the general $Q$-tensor model is of the form
(1.1) $\left\{\begin{aligned} \partial_{t}u+\left(u\cdot\nabla\right)u+\nabla\pi&=\Delta u+\operatorname{div}\bigl(\tau(Q,H(Q))+\sigma(Q,H(Q))\bigr),\kern 5.0pt\operatorname{div}u=0,\\ \partial_{t}Q+\left(u\cdot\nabla\right)Q-S(\nabla u,Q)&=H(Q).\end{aligned}\right$.
Here $H(Q)=\Delta Q-Q+(Q^{2}-\operatorname{tr}(Q^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}})-\operatorname{tr}(Q^{2})Q$ is the variational derivative of the free energy functional. The non-Newtonian stresses due to the molecular orientation of the biaxial liquid crystals
$\tau(Q,H(Q))$ $=-\nabla Q\odot\nabla Q-\xi(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})H(Q)-\xi H(Q)(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})+2\xi(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}\bigl(QH(Q)\bigr)\kern 5.0pt\text{and}\kern 5.0pt$
$\sigma(Q,H(Q))$ $=QH(Q)-H(Q)Q=Q\Delta Q-\Delta QQ$
correspond to the symmetric and antisymmetric part of the stress tensor. For the deformation tensor $\mathbb{D}(u)$ and its antisymmetric counterpart $\mathbb{W}(u)$, the term $S(\nabla u,Q)$ accounting for stretching and rotation is given by $S(\nabla u,Q)=(\xi\mathbb{D}(u)+\mathbb{W}(u))(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})+(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})(\xi\mathbb{D}(u)-\mathbb{W}(u))-2\xi(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(Q\nabla u)$.
Beris-Edwards models for liquid crystals have been investigated by many authors considering simplifications or modifications of the model. In the context of weak solutions, we refer for instance to the articles [6, 33, 34, 21, 16, 1, 47]. Strong solutions were e. g. studied in [2, 11, 38, 30]. These results either concern the 2D case, the situation of $\xi=0$ or simplified variants of the model. Only very recently, a rather complete understanding of the strong well-posedness and dynamics of the general Beris-Edwards model with arbitrary ratio of tumbling and alignment effects $\xi\in\mathbb{R}$ in 3D was obtained in [22].
In the present article, we build on this new development and extend the analysis to the associated interaction problem with a rigid body. The physical motivation for liquid crystal interaction problems is the formation of liquid crystal colloids by dispersion of colloidal particles in the liquid crystal host medium, where we view colloidal particles as rigid bodies. We refer for instance to [31, 29] for more details on the underlying physics. The interaction with a rigid body poses significant challenges. Most importantly, the resulting problem is a moving domain problem, and the equations describing the dynamics of the liquid crystals are coupled to the equations governing the motion of the rigid body via the complicated stress tensor and the continuity of the fluid and rigid body velocity on their moving interface.
More precisely, the dynamics of the rigid body follows Newton’s laws. Consequently, for the total stress tensor $\Sigma(u,\pi,Q)=2\mathbb{D}(u)+\tau(Q,H(Q))+\sigma(Q,H(Q))-\pi\operatorname{Id}_{3}$, the mass $m_{\mathcal{S}}$ and inertia matrix $J$ of the rigid body, the position of its center of mass $h$, its angular velocity $\Omega$, its domain $\mathcal{S}(t)$ and the outward unit normal $\nu(t)$, the rigid body motion is determined by
(1.2) $m_{\mathcal{S}}h^{\prime\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}\Sigma(u,\pi,Q)\nu(t)\,\mathrm{d}\Gamma\kern 5.0pt\text{and}\kern 5.0pt(J\Omega)^{\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}(x-h(t))\times\Sigma(u,\pi,Q)\nu(t)\,\mathrm{d}\Gamma.$
In particular, the complex effects due to the non-Newtonian stresses $\tau$ and $\sigma$ enter in these equations, and estimating the boundary integrals leads to difficulties. The domain of the present problem changes in time due to the motion of the rigid body. The other coupling of the dynamics of the body to the $Q$-tensor model arises from the equality of fluid and body velocity on their interface $\partial\mathcal{S}(t)$ and reads as
(1.3) $u(t,x)=h^{\prime}(t)+\Omega(t)\times(x-h(t)),\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\kern 5.0ptx\in\partial\mathcal{S}(t).$
For handling the present moving boundary problem given in simplified form in (1.1)–(1.3), we first employ a local transformation to a fixed domain. The resulting problem is a quasilinear non-autonomous one, and we tackle it by establishing maximal $\mathrm{L}^{p}$-regularity for a suitable linearization first.
The $Q$-tensor model (1.1) is a mixed-order system, where the diagonal parts are of second order, while the off-diagonal parts are of order one and three. In contrast to the rather well-understood case of systems with principal part, mixed-order systems lack a unified theory. Here we choose an anisotropic ground space for the liquid crystal part, so every entry is of highest order. Maximal regularity is only known to be valid on $\mathrm{L}^{2}\times\mathrm{H}^{1}$, see [22], or on $\mathrm{L}^{q}\times\mathrm{W}^{1,q}$ for $q$ close to $2$, see [3, Section 7.6]. It is an interesting open question whether maximal regularity for the linearized general $Q$-tensor model can be achieved for general $q\in(1,\infty)$ as well. Previous works in this direction are limited to simplified models.
The block structure with all blocks of highest order also leads to significant challenges. We employ here a monolithic approach, so we keep the interface condition (1.3) in the study of the linearized problem and solve the linearization of (1.1) and (1.2) simultaneously. The added mass operator together with an involved lifting procedure allows us to write the linearized problem in terms of a so-called fluid-structure operator tailored to the present setting of the $Q$-tensor model.
We establish the maximal $\mathrm{L}^{p}$-regularity on the anisotropic space $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}$, where $\mathcal{F}_{0}$ is the initial fluid domain, and $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ denotes the space of solenoidal vector fields in $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. In particular, the condition $Q\in\mathbb{S}_{0}^{3}$ is included in the function spaces in elegant way. The maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity with exponential decay for the linearization at zero follows from spectral theory.
In order to obtain the first main result of this paper, the local strong well-posedness of the interaction problem (1.1)–(1.3) for large initial data made precise below in Theorem 4.1, we estimate the nonlinear terms and employ the maximal regularity. The anisotropic ground space $\mathrm{L}^{2}\times\mathrm{H}^{1}$ for the liquid crystals part renders the nonlinear problem more complicated. The time interval of existence in this case is limited by the requirement that we exclude the case of collision of the rigid body with the outer fluid boundary.
With regard to the second main result of the present article, the global strong well-posedness and exponential decay of the solution for small initial data as asserted in Theorem 4.2, we first prove that the total energy of the interaction problem (1.1)–(1.3) is decreasing, which reveals the shape of the equilibria. In a second step, we make use of a fixed point argument, relying on the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of the linearization at zero as well as adjusted decay estimates of the transform. Let us emphasize that we are especially able to show that no collision of the rigid body with the outer fluid boundary occurs. This follows from the exponential stability of the semigroup associated to the linearized problem around zero and the smallness of the initial data. The exponential stability also reveals that all velocities tend to zero as time tends to infinity. However, it remains an interesting open problem to determine if the position of the center of the moving rigid body converges to some point $h_{\infty}\in\mathbb{R}^{3}$ as $t\rightarrow\infty$. In the situation of a rigid body immersed in a viscous, incompressible Newtonian fluid filling the entire $\mathbb{R}^{3}$, this problem has recently been tackled in the case of a ball in [15], then for a rigid body of arbitrary shape in [28] and for the case of several bodies in [10].
Fluid-rigid body interaction problems of incompressible, Newtonian fluids are classical and abundantly documented. We mention e. g. the pioneering article of Serre [39] and the papers [14, 12, 32] in the weak setting or the works [17, 18, 43, 19, 15] in the strong setting, see also [25] for an overview of the topic. Most of the above articles deal with the semilinear setting, whereas here, due to the quasilinear structure of the stress tensor, we need to deal with a quasilinear fluid-structure interaction problem.
In contrast to the case of Newtonian fluids, the investigation of fluid-rigid body interaction problems of liquid crystals started only recently. In [20], Geng, Roy and Zarnescu showed the global existence of a weak solution to the interaction problem of a simplified $Q$-tensor model, whereas here, we study the general $Q$-tensor model with arbitrary ratio $\xi\in\mathbb{R}$ and especially obtain uniqueness of solutions. In [9], the local strong well-posedness and the global strong well-posedness for initial data close to equilibria of the fluid-rigid body interaction problem of the simplified Ericksen-Leslie model (see [26]) were proved. The complexity of the stress tensor, the mixed-order character of the underlying $Q$-tensor model as well as the fact that the molecular orientation is described by a matrix render the analysis of the present interaction problem fundamentally different.
For convenience of the reader, let us give a brief, non-technical overview of the procedure to obtain the well-posedness results of the present interaction problem of a rigid body with the general Beris-Edwards $Q$-tensor model. First, we determine the set of equilibrium solutions by examining the critical points of the energy functional associated with the interaction problem. Next, we employ a transform that only acts locally, i. e., in a neighborhood of the translating and rotating rigid body. By continuity of the body velocities, at least for short time, we can guarantee that there is no collision of the rigid body with the outer fluid boundary. After transforming the moving boundary problem to a fixed domain, we perform the analysis of the linearized problem. The key idea here is to use a lifting procedure in order to eliminate the pressure from the surface integrals of the stress tensor, and to reformulate the linearized problem in operator form. This is done by means of the so-called added mass operator. The idea of this concept can be traced back to Bessel, see [41, 42], and has become classical in the investigation of fluid-rigid body interaction problems. Thanks to a refined spectral analysis, we prove the exponential stability for the linearization at zero, which, together with the smallness of the initial data, helps to show that no collision of the rigid body with the outer fluid boundary occurs for sufficiently small initial data. The proofs of the well-posedness results are based on a fixed point procedure, relying in turn on the analysis of the linearized problem, and they are complemented by suitable nonlinear estimates, where we efficiently handle the transform. Let us emphasize that the complicated nature of the stress tensor, in particular in the terms with the molecular orientation $Q$, leads to involved terms that need to be estimated.
This article has the following outline. Section 2 is dedicated to the introduction of the Beris-Edwards model and the interaction problem with a rigid body. The purpose of Section 3 is to show that the total energy is non-increasing, which also enables us to determine the set of equilibria. We then assert the main results of the paper, Theorem 4.1 on the local strong well-posedness for large data and Theorem 4.2 on the global strong well-posedness and exponential decay for small data, in Section 4. In Section 5, we transform the interaction problem to a fixed domain. Section 6 is devoted to the linear theory; we reformulate the linearized problem by introducing an associated fluid-structure operator and then verify its maximal $\mathrm{L}^{p}$-regularity. The maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity and exponential decay for the linearization around zero follow from spectral analysis. In Section 7, we prove Theorem 4.1 based on the maximal regularity and nonlinear estimates. Section 8 finally discusses the proof of Theorem 4.2.
2. The fluid-rigid body interaction problem of the general $Q$-tensor model
First, we describe the general $Q$-tensor model. By $\mathcal{F}_{T}$, we denote the spacetime associated to the fluid domain and made precise below. The variables are the fluid velocity $u^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{R}^{3}$, the pressure $\pi^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{R}$ and the molecular orientation of the liquid crystal $Q^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{S}_{0,\mathbb{R}}^{3}$. Here $\mathbb{S}_{0,\mathbb{R}}^{3}$ denotes the space of symmetric $3\times 3$-matrices with trace zero, so $\mathbb{S}_{0,\mathbb{R}}^{3}\coloneqq\{M\in\mathbb{R}^{3\times 3}:M=M^{\top}\kern 5.0pt\text{and}\kern 5.0pt\operatorname{tr}M=0\}$. The complex version $\mathbb{S}_{0,\mathbb{C}}^{3}$ is defined analogously. We also write $\mathbb{S}_{0}^{3}$ if no confusion arises. For $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, we denote by $\mathbb{M}_{0,\mathbb{K}}^{3}\coloneqq\{M\in\mathbb{K}^{3\times 3}:\operatorname{tr}M=0\}$ the space of traceless matrices in $\mathbb{K}^{3\times 3}$ and also occasionally omit the subscript K. More background on $Q$-tensor models, including the relation with the Ericksen-Leslie model, can be found in [6], see also [22, Section 2].
One feature of the $Q$-tensor model is the free energy functional taking the form
(2.1) $\mathrm{E}_{Q}(Q^{\mathcal{F}})=\int_{\mathcal{F}(t)}\Bigl(\frac{\lambda}{2}|\nabla Q^{\mathcal{F}}(x)|^{2}+\frac{a}{2}\operatorname{tr}\bigl((Q^{\mathcal{F}}(x))^{2}\bigr)-\frac{b}{3}\operatorname{tr}\bigl((Q^{\mathcal{F}}(x))^{3}\bigr)\Bigr)+\frac{c}{4}\operatorname{tr}\bigl((Q^{\mathcal{F}}(x))^{2}\bigr)^{2}\,\mathrm{d}x$
for constants $a$, $\lambda>0$ and $b$, $c\in\mathbb{R}$. The first addend in (2.1) corresponds to the elastic energy, while the other terms are related to the Landau-de Gennes thermotropic energy, see [13].
For $\mathbb{D}\coloneqq\mathbb{D}(u^{\mathcal{F}})=\nicefrac{{1}}{{2}}(\nabla u^{\mathcal{F}}+(\nabla u^{\mathcal{F}})^{\top})$, and in the presence of tumbling and alignment effects, the full stress tensor with non-Newtonian stresses reads as
(2.2) $\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})=2\mu\mathbb{D}(u^{\mathcal{F}})+\tau(Q^{\mathcal{F}},H(Q^{\mathcal{F}}))+\sigma(Q^{\mathcal{F}},H(Q^{\mathcal{F}}))-\pi^{\mathcal{F}}\operatorname{Id}_{3},$
where $\mu>0$ represents a viscosity parameter. Using the one constant approximation for the Oseen-Frank energy of liquid crystals and a Landau-de Gennes expression for the bulk energy, the variational derivative $H(Q^{\mathcal{F}})$ of the free energy functional $\mathrm{E}_{Q}$ from (2.1) is
(2.3) $H(Q^{\mathcal{F}})\coloneqq\lambda\Delta Q^{\mathcal{F}}-aQ^{\mathcal{F}}+b\bigl((Q^{\mathcal{F}})^{2}-\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)\nicefrac{{\operatorname{Id}_{3}}}{{3}}\bigr)-c\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)Q^{\mathcal{F}},$
while the symmetric and antisymmetric part of the stress tensor are given by
(2.4) $\tau(Q^{\mathcal{F}})=\tau(Q^{\mathcal{F}},H(Q^{\mathcal{F}}))$ $\coloneqq-\lambda\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}}-\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})H(Q^{\mathcal{F}})-\xi H(Q^{\mathcal{F}})(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})$
$\quad+2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}\bigl(Q^{\mathcal{F}}H(Q^{\mathcal{F}})\bigr)\kern 5.0pt\text{and}\kern 5.0pt$
$\sigma(Q^{\mathcal{F}})=\sigma(Q^{\mathcal{F}},H(Q^{\mathcal{F}}))$ $\coloneqq Q^{\mathcal{F}}H(Q^{\mathcal{F}})-H(Q^{\mathcal{F}})Q^{\mathcal{F}}=\lambda\bigl(Q^{\mathcal{F}}\Delta Q^{\mathcal{F}}-\Delta Q^{\mathcal{F}}Q^{\mathcal{F}}\bigr),$
respectively. Here the parameter $\xi\in\mathbb{R}$ measures the ratio of tumbling and alignment effects, and the $(i,j)$-component of $\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}}$ equals $\operatorname{tr}(\partial_{i}Q^{\mathcal{F}}\partial_{j}Q^{\mathcal{F}})$. With $\mathbb{W}\coloneqq\mathbb{W}(u^{\mathcal{F}})=\nicefrac{{1}}{{2}}(\nabla u^{\mathcal{F}}-(\nabla u^{\mathcal{F}})^{\top})$, we introduce the expression $S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})$ providing information on the way the gradient of the velocity $\nabla u^{\mathcal{F}}$ stretches and rotates the order parameter $Q^{\mathcal{F}}$ and given by
$S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})\coloneqq(\xi\mathbb{D}+\mathbb{W})(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})+(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})(\xi\mathbb{D}-\mathbb{W})-2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}\bigl(Q^{\mathcal{F}}\nabla u^{\mathcal{F}}\bigr).$
For simplicity, we assume the parameters to be equal to $1$ in the sequel, so $\mu=\lambda=a=b=c=1$, because it does not affect the analysis. The resulting Beris-Edwards model is given by
(2.5) $\left\{\begin{aligned} {\partial_{t}u^{\mathcal{F}}}+\left(u^{\mathcal{F}}\cdot\nabla\right)u^{\mathcal{F}}+\nabla\pi^{\mathcal{F}}&=\Delta u^{\mathcal{F}}+\operatorname{div}\bigl(\tau(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})\bigr),\kern 5.0pt\operatorname{div}u^{\mathcal{F}}=0,&&\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{T},\\ \partial_{t}Q^{\mathcal{F}}+\left(u^{\mathcal{F}}\cdot\nabla\right)Q^{\mathcal{F}}-S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})&=H(Q^{\mathcal{F}}),&&\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{T}.\end{aligned}\right$.
In the sequel, we denote by $\mathcal{O}\subset\mathbb{R}^{3}$ a bounded domain with boundary of class $\mathrm{C}^{3}$, and $0<T\leq\infty$ represents a positive time. We consider a rigid body moving in the interior of $\mathcal{O}$, and the body is assumed to be closed, bounded and simply connected. Furthermore, for $t\in(0,T)$, we denote by $\mathcal{S}(t)$ the domain occupied by the rigid body at time $t$. The rest of the domain $\mathcal{F}(t)\coloneqq\mathcal{O}\setminus\mathcal{S}(t)$ is assumed to be filled by a viscous, incompressible fluid with liquid crystals. We use the notation $\mathcal{S}_{0}$ for the initial domain of the rigid body and suppose its boundary $\partial\mathcal{S}_{0}$ to be of class $\mathrm{C}^{3}$ as well. As a result, $\mathcal{F}_{0}=\mathcal{O}\setminus\mathcal{S}_{0}$ is the initial fluid domain. On the other hand, the solid domain $\mathcal{S}(t)$ at time $t$ is given by $\mathcal{S}(t)=\left\{h(t)+\mathbb{O}(t)y:y\in\mathcal{S}_{0}\right\}$, where $h(t)$ is the center of mass of the body, and $\mathbb{O}(t)\in\mathrm{SO}(3)$ is linked to the rotation of the rigid body. Since $\mathbb{O}^{\prime}(t)\mathbb{O}^{\top}(t)$ is skew symmetric, there exists a unique $\Omega(t)\in\mathbb{R}^{3}$ with $\mathbb{O}^{\prime}(t)\mathbb{O}^{\top}(t)y=\Omega(t)$ for all $t>0$ and $y\in\mathbb{R}^{3}$. By $\mathcal{F}_{T}$, we denote the spacetime related to $\mathcal{F}(t)$, i. e., $\mathcal{F}_{T}\coloneqq\{(t,x):t\in(0,T),\,x\in\mathcal{F}(t)\}$, and $\mathcal{S}_{T}$ and $\partial\mathcal{S}_{T}$ corresponding to $\mathcal{S}(t)$ and $\partial\mathcal{S}(t)$ are defined likewise. The rigid body velocity is
$u^{\mathcal{S}}(t,x)=h^{\prime}(t)+\Omega(t)\times(x-h(t)),\kern 5.0pt\text{for all}\kern 5.0pt(t,x)\in\mathcal{S}_{T},$
with the translational velocity $h^{\prime}\colon(0,T)\to\mathbb{R}^{3}$ and the angular velocity of the body $\Omega\colon(0,T)\to\mathbb{R}^{3}$.
We denote by $\rho_{\mathcal{S}}$ the density of the rigid body and assume $\rho_{\mathcal{S}}\equiv 1$. As a result, $m_{\mathcal{S}}=\int_{\mathcal{S}(t)}\,\mathrm{d}x$ is its total mass. Without loss of generality, we suppose the initial position of the center of mass of the body to be at the origin, so $h(0)=0$. Moreover, the inertia matrix $J(t)$ and its initial value $J_{0}$ take the shape
$J_{0}=\int\limits_{\mathcal{S}_{0}}\left(|x|^{2}\operatorname{Id}_{3}-x\otimes x\right)\,\mathrm{d}x\kern 5.0pt\text{as well as}\kern 5.0ptJ(t)=\int\limits_{\mathcal{S}(t)}\left(|x-h(t)|^{2}\operatorname{Id}_{3}-(x-h(t))\otimes(x-h(t))\right)\,\mathrm{d}x.$
Therefore, $J(t)$ is a symmetric, positive definite matrix and fulfills the identities
(2.6) $J(t)=\mathbb{O}(t)J_{0}\mathbb{O}^{\top}(t)\kern 5.0pt\text{and}\kern 5.0ptJ(t)x_{1}\cdot x_{2}=\int\limits_{\mathcal{S}_{0}}(x_{1}\times\mathbb{O}(t)y)\cdot(x_{2}\times\mathbb{O}(t)y)\,\mathrm{d}y,\kern 5.0pt\text{for all}\kern 5.0ptx_{1},x_{2}\in\mathbb{R}^{3}.$
Using the stress tensor $\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})$ from (2.2), and noting that $\nu(t)$ denotes the unit outward normal to the boundary of $\mathcal{F}(t)$, i. e., $\nu(t)$ directed inwards to $\mathcal{S}(t)$, the motion of the rigid body is governed by Newton’s law
(2.7) $m_{\mathcal{S}}h^{\prime\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\,\mathrm{d}\Gamma,\kern 5.0pt(J\Omega)^{\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}(x-h(t))\times\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\,\mathrm{d}\Gamma.$
Next, we discuss the boundary and interface conditions. We assume the fluid velocity to satisfy homogeneous boundary conditions on the outer boundary and to coincide with the rigid body velocity on their common interface, whereas $Q^{\mathcal{F}}$ fulfills homogeneous Neumann boundary conditions, i. e.,
(2.8) $u^{\mathcal{F}}=0,\kern 5.0pt\partial_{\nu}Q^{\mathcal{F}}=0,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0pt\text{and}\kern 5.0ptu^{\mathcal{F}}=u^{\mathcal{S}},\kern 5.0pt\partial_{\nu}Q^{\mathcal{F}}=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{T}.$
The last ingredient for the full system are the initial conditions
(2.9) $u^{\mathcal{F}}(0,x)=v_{0}(x),\kern 5.0ptQ^{\mathcal{F}}(0,x)=Q_{0}(x),\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{0},\kern 5.0pth(0)=0,\kern 5.0pth^{\prime}(0)=\ell_{0}\kern 5.0pt\text{and}\kern 5.0pt\Omega(0)=\omega_{0}.$
Let us conclude this section by settling some more concepts required in this article. First, we make precise function spaces on time-dependent domains. For this purpose, let $X\colon\mathcal{F}_{T}\to\mathcal{F}_{0}$ be a map with $\varphi\colon\mathcal{F}_{T}\to(0,T)\times\mathcal{F}_{0}$, $(t,x)\mapsto(t,X(t,x))$ being a $\mathrm{C}^{1}$-diffeomorphism, and $X(\tau,\cdot)\colon\mathcal{F}(\tau)\to\mathcal{F}_{0}$ being $\mathrm{C}^{3}$-diffeomorphisms for all $\tau\in[0,T]$. For $p\in(1,\infty)$, $s\in\{0,1\}$ and $l\in\{0,1,2,3\}$, we then set
$\mathrm{W}^{s,p}\bigl(0,T;\mathrm{H}^{l}(\mathcal{F}(\cdot))\bigr)\coloneqq\bigl\{f(t,\cdot)\colon\mathcal{F}(t)\to\mathbb{R}:f\circ\varphi\in\mathrm{W}^{s,p}\bigl(0,T;\mathrm{H}^{l}(\mathcal{F}_{0})\bigr)\bigr\}.$
In this paper, $\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ represents the $\mathrm{L}^{2}(\mathcal{F}_{0})$-functions with spatial average zero, while $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ denotes the solenoidal vector fields on $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$, i. e., for the Helmholtz projection $\mathbb{P}$, we define $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\coloneqq\mathbb{P}\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. For $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})$, $\mathrm{Y}_{\gamma}\coloneqq\mathrm{B}_{2p}^{2-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0})^{3}\cap\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}$ and $p>\nicefrac{{4}}{{3}}$, we define
(2.10) $\mathrm{X}_{\gamma}\coloneqq\{z_{0}\in\mathrm{Y}_{\gamma}:v_{0}=0,\kern 5.0pt\partial_{n}Q_{0}=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{O},\kern 5.0ptv_{0}=\ell_{0}+\omega_{0}\times y,\kern 5.0pt\partial_{n}Q_{0}=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{0}\},$
where $\mathrm{B}_{2p}^{s}(\mathcal{F}_{0})$ denotes the Besov space. The traces and normal derivatives are well-defined due to $p>\nicefrac{{4}}{{3}}$, see e. g. [44, Theorem 4.7.1]. In the main results, we will assume that $p>4$ in order to establish suitable estimates of the nonlinear terms. These estimates, carried out in detail in Section 7 and Section 8, are also based on Sobolev embeddings, which require that $p>4$. If $p\in(1,\nicefrac{{4}}{{3}})$, then we set $\mathrm{X}_{\gamma}\coloneqq\{z_{0}\in\mathrm{Y}_{\gamma}:v_{0}\cdot n=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{O},\kern 5.0ptv_{0}\cdot n=(\ell_{0}+\omega_{0}\times y)\cdot n,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{0}\}$, where $n$ denotes the outer unit normal vector to the fluid boundary, so it is directed towards $\partial\mathcal{S}_{0}$.
3. Energy decay and equilibria of the interaction problem
In this section, we show that the total energy of the interaction problem (2.5)–(2.9) is a non-increasing functional, and we deduce the set of equilibria. First, we decompose $\tau=\tau_{\mathrm{h}}+\tau_{\mathrm{l}}$ and $H=H_{\mathrm{h}}+H_{\mathrm{l}}$ into the respective higher and lower order terms. Invoking the commutator [A,B]=ABBA[A,B]=AB-BA and anticommutator $\{A,B\}=AB+BA$ of $A$, $B\in\mathbb{C}^{3\times 3}$, and writing $\mathbb{D}=\mathbb{D}(u^{\mathcal{F}})$ and $\mathbb{W}=\mathbb{W}(u^{\mathcal{F}})$, we get
(3.1) $S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})$ $=-[Q^{\mathcal{F}},\mathbb{W}]+\xi\bigl(\nicefrac{{2}}{{3}}\mathbb{D}+\{Q^{\mathcal{F}},\mathbb{D}\}-2(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}\bigl(Q^{\mathcal{F}}\nabla u^{\mathcal{F}}\bigr)\bigr),$
$H_{\mathrm{h}}(Q^{\mathcal{F}})$ $=\Delta Q^{\mathcal{F}}-Q^{\mathcal{F}},\kern 5.0ptH_{\mathrm{l}}(Q^{\mathcal{F}})=\bigl((Q^{\mathcal{F}})^{2}-\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)\nicefrac{{\operatorname{Id}_{3}}}{{3}}\bigr)-\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)Q^{\mathcal{F}},$
$\sigma(Q^{\mathcal{F}})$ $=-[Q^{\mathcal{F}},(-\Delta+\operatorname{Id}_{3})Q^{\mathcal{F}}],$
$\tau_{\mathrm{h}}(Q^{\mathcal{F}})$ $=-\xi\bigl(\nicefrac{{2}}{{3}}H_{\mathrm{h}}+\{Q^{\mathcal{F}},H_{\mathrm{h}}\}-2(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(Q^{\mathcal{F}}H_{\mathrm{h}})\bigr)\kern 5.0pt\text{and}\kern 5.0pt$
$\tau_{\mathrm{l}}(Q^{\mathcal{F}})$ $=2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\bigl(\operatorname{tr}\bigl((Q^{\mathcal{F}})^{3}\bigr)-\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)^{2}\bigr)-\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}}-2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})H_{\mathrm{l}}.$
For $Q^{\mathcal{F}}\in\mathbb{S}_{0,\mathbb{C}}^{3}$, we define the maps $S_{\xi}(Q^{\mathcal{F}})\in\mathcal{L}(\mathbb{S}_{0,\mathbb{C}}^{3},\mathbb{M}_{0,\mathbb{C}}^{3})$ as well as $\tilde{S}_{\xi}(Q^{\mathcal{F}})\in\mathcal{L}(\mathbb{M}_{0,\mathbb{C}}^{3},\mathbb{S}_{0,\mathbb{C}}^{3})$ by
(3.2) $S_{\xi}(Q^{\mathcal{F}})A$ $\coloneqq[Q^{\mathcal{F}},A]-\nicefrac{{2\xi}}{{3}}A-\xi\{Q^{\mathcal{F}},A\}+2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(Q^{\mathcal{F}}A)\kern 5.0pt\text{and}\kern 5.0pt$
$\tilde{S}_{\xi}(Q^{\mathcal{F}})B$ $\coloneqq[\overline{Q^{\mathcal{F}}},\nicefrac{{1}}{{2}}(B-B^{\top})]-\nicefrac{{\xi}}{{3}}(B+B^{\top})-\xi\{\overline{Q^{\mathcal{F}}},\nicefrac{{1}}{{2}}(B+B^{\top})\}$
$\quad+\xi(\overline{Q^{\mathcal{F}}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(\overline{Q^{\mathcal{F}}}(B+B^{\top})).$
For real $u^{\mathcal{F}}$, $Q^{\mathcal{F}}$, we get $\tau_{\mathrm{h}}(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})=S_{\xi}(Q^{\mathcal{F}})(\Delta-\operatorname{Id}_{3})Q^{\mathcal{F}}$ and $S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})=\tilde{S}_{\xi}(Q^{\mathcal{F}})\nabla u^{\mathcal{F}}$. The following lemma, for which we refer to [22, Lemma 5.1], establishes a link between $S_{\xi}$ and $\tilde{S}_{\xi}$. Note that we also denote the product $\langle A,B\rangle_{\mathbb{C}^{3\times 3}}$ by $A:B$, and it is defined by $\operatorname{tr}(A\overline{B})$.
Lemma 3.1.
For $\xi\in\mathbb{R}$, $Q^{\mathcal{F}}\in\mathbb{S}_{0}^{3}$, $A\in\mathbb{S}_{0}^{3}$ and $B\in\mathbb{M}_{0}^{3}$, we have $\langle S_{\xi}(Q^{\mathcal{F}})A,B\rangle_{\mathbb{C}^{3\times 3}}=\langle A,\tilde{S}_{\xi}(Q^{\mathcal{F}})B\rangle_{\mathbb{C}^{3\times 3}}$.
Now, we investigate the total energy of the interaction problem (2.5)–(2.9). Recalling the free energy functional $\mathrm{E}_{Q}$ from (2.1), we define the total energy, which additionally comprises the kinetic energy and the energy resulting from the rigid body motion, by
(3.3) $\mathrm{E}(t)\coloneqq\mathrm{E}_{\mathrm{kin}}(t)+\mathrm{E}_{Q}(t)+\mathrm{E}_{\mathrm{trans}}(t)+\mathrm{E}_{\mathrm{rot}}(t),\kern 5.0pt\text{where}\kern 5.0pt$
$\mathrm{E}_{\mathrm{kin}}(t)=\frac{1}{2}\int_{\mathcal{F}(t)}|u^{\mathcal{F}}(t,x)|^{2}\,\mathrm{d}x,\kern 5.0pt\mathrm{E}_{\mathrm{trans}}(t)=\frac{1}{2}m_{\mathcal{S}}|h^{\prime}(t)|^{2}\kern 5.0pt\text{and}\kern 5.0pt\mathrm{E}_{\mathrm{rot}}(t)=\frac{1}{2}(J\Omega\cdot\Omega)(t).$
The next proposition discusses $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathrm{E}(t)$ and the resulting set of equilibria $\mathcal{E}$ to (2.5)–(2.8).
Proposition 3.2.
Let $p>4$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, for $\mathrm{X}_{\gamma}$ as introduced in (2.10), recall $H$ from (2.3), and consider a solution $(u^{\mathcal{F}},Q^{\mathcal{F}},h^{\prime},\Omega,\pi^{\mathcal{F}})$, whose existence will be discussed in Theorem 4.1, to (2.5)–(2.9) in the regularity class
$\mathrm{W}^{1,p}\bigl(0,T;\mathrm{L}^{2}(\mathcal{F}(\cdot))^{3}\bigr)\cap\mathrm{L}^{p}\bigl(0,T;\mathrm{H}^{2}(\mathcal{F}(\cdot))^{3}\bigr)\times\mathrm{W}^{1,p}\bigl(0,T;\mathrm{H}^{1}(\mathcal{F}(\cdot),\mathbb{S}_{0,\mathbb{R}}^{3})\bigr)\cap\mathrm{L}^{p}\bigl(0,T;\mathrm{H}^{3}(\mathcal{F}(\cdot),\mathbb{S}_{0,\mathbb{R}}^{3})\bigr)$
$\times\mathrm{W}^{1,p}(0,T)^{3}\times\mathrm{W}^{1,p}(0,T)^{3}\times\mathrm{L}^{p}\bigl(0,T;\mathrm{H}^{1}(\mathcal{F}(\cdot))\cap\mathrm{L}_{0}^{2}(\mathcal{F}(\cdot))\bigr).$
  1. (a)
    The energy functional $\mathrm{E}(t)$ from (3.3) satisfies $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}(t)=-\int_{\mathcal{F}(t)}|\nabla u^{\mathcal{F}}|^{2}\,\mathrm{d}x-\int_{\mathcal{F}(t)}\operatorname{tr}(H^{2})\,\mathrm{d}x\leq 0$.
  2. (b)
    We have $\mathcal{E}=\{(0,Q_{*},0,0):H(Q_{*})=0,\text{ in }\mathcal{F}_{0},\kern 5.0pt\partial_{\nu}Q_{*}=0,\text{ on }\partial\mathcal{F}_{0}\}$ for the set of equilibria.
Proof.
Mimicking the procedure from the proof of [9, Proposition 3.2] with the stress tensor $\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})$ from (2.2), using integration by parts in conjunction with the Reynolds transport theorem and the boundary conditions (2.8), for $\tau$ and $\sigma$ from (2.4), and with the outer unit normal vector $n$, we infer that
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{kin}}(t)=-\int_{\mathcal{F}(t)}|\nabla u^{\mathcal{F}}|^{2}\,\mathrm{d}x-\int_{\mathcal{F}(t)}(\tau(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})):\nabla u^{\mathcal{F}}\,\mathrm{d}x-\int_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\cdot u^{\mathcal{F}}\,\mathrm{d}\Gamma.$
Similarly as in [9, (3.15) and (3.17)], we find $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{trans}}(t)+\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{rot}}(t)=\int_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\cdot u^{\mathcal{F}}\,\mathrm{d}\Gamma$. A concatenation of the latter two identities leads to
(3.4) $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{kin}}(t)+\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{trans}}(t)+\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{rot}}(t)=-\int_{\mathcal{F}(t)}|\nabla u^{\mathcal{F}}|^{2}\,\mathrm{d}x-\int_{\mathcal{F}(t)}(\tau(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})):\nabla u^{\mathcal{F}}\,\mathrm{d}x.$
It remains to handle $\mathrm{E}_{Q}$ from (2.1). For this purpose, we multiply the equation (2.5)2 satisfied by the molecular orientation $Q^{\mathcal{F}}$ from the right by H-H and integrate over $\mathcal{F}(t)$. This yields
(3.5) $\int_{\mathcal{F}(t)}\partial_{t}Q^{\mathcal{F}}:H\,\mathrm{d}x-\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:H\,\mathrm{d}x+\int_{\mathcal{F}(t)}S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}}):H\,\mathrm{d}x=-\int_{\mathcal{F}(t)}\operatorname{tr}(H^{2})\,\mathrm{d}x.$
From Section 3 and the aforementioned identities $\tau_{\mathrm{h}}(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})=S_{\xi}(Q^{\mathcal{F}})((\Delta-\operatorname{Id}_{3})Q^{\mathcal{F}})$ as well as $S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})=\tilde{S}_{\xi}(Q^{\mathcal{F}})\nabla u^{\mathcal{F}}$ and $S_{\xi}(Q^{\mathcal{F}})H_{\mathrm{l}}=\tau_{\mathrm{l}}(Q^{\mathcal{F}})+\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}}$, we deduce that
(3.6) $\int_{\mathcal{F}(t)}S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}}):H\,\mathrm{d}x=-\int_{\mathcal{F}(t)}\nabla u^{\mathcal{F}}:(\tau(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}}))\,\mathrm{d}x-\int_{\mathcal{F}(t)}\nabla u^{\mathcal{F}}:(\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}})\,\mathrm{d}x.$
Next, we calculate $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathrm{E}_{Q}$. In fact, the Reynolds transport theorem and computations, also relying on the identity $\operatorname{tr}(\partial_{t}Q^{\mathcal{F}})=\partial_{t}\operatorname{tr}(Q^{\mathcal{F}})=0$ by $Q^{\mathcal{F}}\in\mathbb{S}_{0,\mathbb{R}}^{3}$, yield that $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{Q}(t)$ is given by
$-\int_{\mathcal{F}(t)}\partial_{t}Q^{\mathcal{F}}:H\,\mathrm{d}x+\int_{\partial\mathcal{S}(t)}\Bigl(\frac{1}{2}|\nabla Q^{\mathcal{F}}|^{2}+\frac{1}{2}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)-\frac{1}{3}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{3}\bigr)+\frac{1}{4}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)^{2}\Bigr)(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma.$
Putting this together with (3.4), and plugging in (3.5) as well as (3.6), we conclude that $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}(t)$ equals
(3.7) $-\int_{\mathcal{F}(t)}|\nabla u^{\mathcal{F}}|^{2}\,\mathrm{d}x-\int_{\mathcal{F}(t)}\operatorname{tr}(H^{2})\,\mathrm{d}x+\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:H\,\mathrm{d}x+\int_{\mathcal{F}(t)}\nabla u^{\mathcal{F}}:(\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}})\,\mathrm{d}x$
$\int_{\partial\mathcal{S}(t)}\Bigl(\frac{1}{2}|\nabla Q^{\mathcal{F}}|^{2}+\frac{1}{2}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)-\frac{1}{3}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{3}\bigr)+\frac{1}{4}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)^{2}\Bigr)(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma.$
For verifying (a), it thus remains to show that the last three terms in (3.7) cancel out. To this end, we address $\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:H\,\mathrm{d}x$ and discuss the resulting integrals for each term in $H$. Integrating by parts, and employing $\operatorname{div}u^{\mathcal{F}}=0$, (2.8) and the identities $(u^{\mathcal{F}}\cdot\nabla)((Q^{\mathcal{F}})^{2})=(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}\cdot Q^{\mathcal{F}}+Q^{\mathcal{F}}\cdot(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}$, $\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:(\operatorname{tr}((Q^{\mathcal{F}})^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}})\,\mathrm{d}x=0$ thanks to $\operatorname{tr}(Q^{\mathcal{F}})=0$ as well as the identity involving the above inner product $Q^{\mathcal{F}}:\bigl((u^{\mathcal{F}}\cdot\nabla)(\operatorname{tr}((Q^{\mathcal{F}})^{2})Q^{\mathcal{F}})\bigr)=3(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:\bigl(\operatorname{tr}((Q^{\mathcal{F}})^{2})Q^{\mathcal{F}}\bigr)$, we find that
(3.8) $\quad\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:\Bigl(-Q^{\mathcal{F}}+(Q^{\mathcal{F}})^{2}-\operatorname{tr}((Q^{\mathcal{F}})^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}}-\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)Q^{\mathcal{F}}\Bigr)\,\mathrm{d}x$
$=-\int_{\partial\mathcal{S}(t)}\Bigl(\frac{1}{2}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)-\frac{1}{3}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{3}\bigr)+\frac{1}{4}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)^{2}\Bigr)(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma.$
By (2.3), it remains to consider the product with $\Delta Q^{\mathcal{F}}$. Integrating by parts, we obtain
$\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:\Delta Q^{\mathcal{F}}\,\mathrm{d}x$ $=-\int_{\mathcal{F}(t)}\nabla u^{\mathcal{F}}:(\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}})\,\mathrm{d}x-\sum_{i,j,k=1}^{3}\int_{\mathcal{F}(t)}u_{k}^{\mathcal{F}}\partial_{k}\nabla Q_{ij}^{\mathcal{F}}\cdot\nabla Q_{ji}^{\mathcal{F}}\,\mathrm{d}x.$
For the second term, we obtain $\sum_{i,j,k=1}^{3}\int_{\mathcal{F}(t)}u_{k}^{\mathcal{F}}\partial_{k}\nabla Q_{ij}^{\mathcal{F}}\cdot\nabla Q_{ji}^{\mathcal{F}}\,\mathrm{d}x=-\frac{1}{2}\int_{\partial\mathcal{S}(t)}|\nabla Q^{\mathcal{F}}|^{2}(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma$, so
(3.9) $\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:\Delta Q^{\mathcal{F}}\,\mathrm{d}x=-\int_{\mathcal{F}(t)}\nabla u^{\mathcal{F}}:(\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}})\,\mathrm{d}x-\frac{1}{2}\int_{\partial\mathcal{S}(t)}|\nabla Q^{\mathcal{F}}|^{2}(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma.$
Inserting (3.8) and (3.9) into (3.7), we recover the assertion of (a).
Concerning (b), we argue that the equilibria $\mathcal{E}$ are the critical points of the energy functional $\mathrm{E}$, so assume $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathrm{E}(t)\right|_{t=t_{0}}=0$ at $t_{0}>0$. From (a), it follows that $\nabla u^{\mathcal{F}}=0$ and thus also $u^{\mathcal{F}}=0$ in $\mathcal{F}(t_{0})$ thanks to $u^{\mathcal{F}}=0$ on $\partial\mathcal{O}$. As in the proof of [37, Lemma 2.1], we deduce that $h^{\prime}=\Omega=0$. The relation in (a) also shows that $H(Q_{*})=0$ has to hold for an equilibrium, so $\mathcal{E}$ is of the asserted shape. ∎
Remark $3.$
Combining Section 3(b) and [22, Remark 3.3], stating that $H(Q_{*})=0$ implies $Q_{*}=0$ for spatially constant $Q_{*}$, we conclude that $(0,0,0,0)$ is the only such equilibrium to (2.5)–(2.8).
4. Main results
The first main result asserts the local strong well-posedness of the interaction problem for initial data in $\mathrm{X}_{\gamma}$. In contrast to [20], which establishes the existence of a weak solution to a slightly simplified version of the $Q$-tensor model, we consider here the general $Q$-tensor model with arbitrary ratio $\xi$ of tumbling and alignment effects in the strong setting, and we obtain uniqueness of the strong solution.
Theorem 4.1.
Let $p>4$, consider a bounded domain $\mathcal{O}\subset\mathbb{R}^{3}$ of class $\mathrm{C}^{3}$ and the domains $\mathcal{S}_{0}$, $\mathcal{F}_{0}$ of the rigid body and the fluid at time zero, and let $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$. If $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$ for some $r>0$, then there exist $T^{\prime}\in(0,T]$ and $X\in\mathrm{C}^{1}([0,T^{\prime}];\mathrm{C}^{2}(\mathbb{R}^{3})^{3})$ so that $X(\tau,\cdot)\colon\mathcal{F}(\tau)\to\mathcal{F}_{0}$ are $\mathrm{C}^{2}$-diffeomorphisms for all $\tau\in[0,T^{\prime}]$, and the interaction problem (2.5)–(2.9) has a unique solution
$u^{\mathcal{F}}$ $\in\mathrm{W}^{1,p}\bigl(0,T^{\prime};\mathrm{L}^{2}(\mathcal{F}(\cdot))^{3}\bigr)\cap\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{2}(\mathcal{F}(\cdot))^{3}\bigr),$
$Q^{\mathcal{F}}$ $\in\mathrm{W}^{1,p}\bigl(0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}(\cdot),\mathbb{S}_{0}^{3})\bigr)\cap\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{3}(\mathcal{F}(\cdot),\mathbb{S}_{0}^{3})\bigr),$
$h^{\prime}$ $\in\mathrm{W}^{1,p}(0,T^{\prime})^{3},\kern 5.0pth\in\mathrm{L}^{\infty}(0,T^{\prime})^{3},\kern 5.0pt\Omega\in\mathrm{W}^{1,p}(0,T^{\prime})^{3}\kern 5.0pt\text{and}\kern 5.0pt\pi^{\mathcal{F}}\in\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}(\cdot))\cap\mathrm{L}_{0}^{2}(\mathcal{F}(\cdot))\bigr).$
A similar result can be obtained with external forces and torques in (2.7) provided they lie in $\mathrm{L}^{p}(0,T)^{3}$. The intersection with $\mathrm{L}_{0}^{2}(\mathcal{F}(\cdot))$ for the pressure rules out constants and thus leads to uniqueness.
It readily follows that the zero solution is an equilibrium solution to (2.5)–(2.8). In Section 3 and Section 3, it has been revealed that the zero solution is the only spatially constant equilibrium. The second main result below states the existence of a unique, global-in-time, strong solution for small data as well as the exponential convergence of this solution to the trivial equilibrium.
Theorem 4.2.
Let $p>4$, let $\mathcal{O}\subset\mathbb{R}^{3}$ be a bounded $\mathrm{C}^{3}$-domain, consider the initial domains of the rigid body and the fluid $\mathcal{S}_{0}$ and $\mathcal{F}_{0}$, and assume $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$, for $r>0$. Then there is $\eta_{0}>0$ so that for all $\eta\in(0,\eta_{0})$, there are constants $\delta_{0}>0$ and $C>0$ such that for all $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\overline{B}_{\mathrm{X}_{\gamma}}(z_{0},\delta)$, with $\delta\in(0,\delta_{0})$, there exists a unique solution $(u^{\mathcal{F}},Q^{\mathcal{F}},h^{\prime},\Omega,\pi^{\mathcal{F}})$ to (2.5)–(2.9) satisfying
$\|\mathrm{e}^{\eta(\cdot)}u^{\mathcal{F}}\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{2}(\mathcal{F}(\cdot)))}+\|\mathrm{e}^{\eta(\cdot)}\partial_{t}u^{\mathcal{F}}\|_{\mathrm{L}^{p}(0,\infty;\mathrm{L}^{2}(\mathcal{F}(\cdot)))}+\|\mathrm{e}^{\eta(\cdot)}u^{\mathcal{F}}\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{B}_{2p}^{2-\nicefrac{{2}}{{p}}}(\mathcal{F}(\cdot)))}$
$+\|\mathrm{e}^{\eta(\cdot)}Q^{\mathcal{F}}\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{3}(\mathcal{F}(\cdot)))}+\|\mathrm{e}^{\eta(\cdot)}\partial_{t}Q^{\mathcal{F}}\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{1}(\mathcal{F}(\cdot)))}+\|\mathrm{e}^{\eta(\cdot)}Q^{\mathcal{F}}\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}(\cdot)))}$
$+\|\mathrm{e}^{\eta(\cdot)}h^{\prime}\|_{\mathrm{W}^{1,p}(0,\infty)}+\|h\|_{\mathrm{L}^{\infty}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\Omega\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\pi^{\mathcal{F}}\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{1}(\mathcal{F}(\cdot))\cap\mathrm{L}_{0}^{2}(\mathcal{F}(\cdot)))}\leq C\delta.$
Furthermore, $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in[0,\infty)$, so no collision with the fluid boundary occurs.
5. Transformation to a fixed domain
This section is devoted to the change of variables to the fixed domain. In the sequel, we present a classical transformation which was e. g. used in [24]. Denoting by $m(t)$ the skew-symmetric matrix satisfying $m(t)x=\Omega(t)\times x$, we investigate
(5.1) $\left\{\begin{aligned} \partial_{t}X_{0}(t,y)&=m(t)(X_{0}(t,y)-h(t))+h^{\prime}(t),&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\mathbb{R}^{3},\\ X_{0}(0,y)&=y,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathbb{R}^{3}.\end{aligned}\right$.
The solution to (5.1) is of the form $X_{0}(t,y)=\mathbb{O}(t)y+h(t)$ for $\mathbb{O}(t)\in\mathrm{SO}(3)$, where $\mathbb{O}\in\mathrm{W}^{2,p}(0,T)^{3\times 3}$ holds true if $h^{\prime}$, $\Omega\in\mathrm{W}^{1,p}(0,T)^{3}$. At the same time, the inverse $Y_{0}(t,x)=\mathbb{O}^{\top}(t)(x-h(t))$ of $X_{0}(t)$ solves $\partial_{t}Y_{0}(t,x)=-M(t)Y_{0}(t,x)-\ell(t)$ on $(0,T)\times\mathbb{R}^{3}$, where $M(t)\coloneqq\mathbb{O}^{\top}(t)m(t)\mathbb{O}(t)$ and $\ell(t)\coloneqq\mathbb{O}^{\top}(t)h^{\prime}(t)$, as well as $Y_{0}(0,x)=x$.
So far, the transformation acts globally in space, so we modify $X_{0}$ and $Y_{0}$ such that it only operates in a suitable open neighborhood of the moving body. For this purpose, we introduce a new diffeomorphism $X$, defined implicitly to be the solution to
(5.2) $\left\{\begin{aligned} \partial_{t}X(t,y)&=b(t,X(t,y)),&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\mathbb{R}^{3},\\ X(0,y)&=y,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathbb{R}^{3}.\end{aligned}\right$.
The right-hand side $b$ will force the transform to fulfill the above requirements. Let us recall the assumption $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$ for some $r>0$. By the continuity of the body velocity, we take into account the solution up to a time for which a small distance such as $\nicefrac{{r}}{{2}}$ of the rigid body and the outer boundary is guaranteed. Thus, we consider $\chi\in\mathrm{C}^{\infty}(\mathbb{R}^{3};[0,1])$ such that $\chi(x)=1$ if $\mathrm{dist}(x,\partial\mathcal{O})\geq r$ and $\chi(x)=0$ if $\mathrm{dist}(x,\partial\mathcal{O})\leq\nicefrac{{r}}{{2}}$. With this cut-off, we set the right-hand side $b\colon[0,T]\times\mathbb{R}^{3}\to\mathbb{R}^{3}$ of (5.2) to be
$b(t,x)\coloneqq\chi(x)[m(t)(x-h(t))+h^{\prime}(t)]-B_{\mathcal{F}_{0}}(\nabla\chi(\cdot)[m(t)(\cdot-h(t))+h^{\prime}(t)])(x),$
with $B_{\mathcal{F}_{0}}\colon\mathrm{C}_{c}^{\infty}(\mathcal{F}_{0})\to\mathrm{C}_{c}^{\infty}(\mathcal{F}_{0})^{3}$ being the Bogovskiĭ operator associated to $\mathcal{F}_{0}$. Similarly as in [9, Section 4], we argue that $b\in\mathrm{W}^{1,p}\bigl(0,T;\mathrm{C}_{c,\sigma}(\mathcal{F}_{0})^{3}\bigr)$, where the subscript σ indicates that $\operatorname{div}(b(t,\cdot))=0$, and $b|_{\partial\mathcal{S}_{0}}=m(x-h)+h^{\prime}$.
Besides, the Picard-Lindelöf theorem yields the existence of a unique solution $X\in\mathrm{C}^{1}\bigl((0,T);\mathrm{C}^{\infty}(\mathbb{R}^{3})^{3}\bigr)$ to (5.2) for given $h^{\prime}$, $\Omega\in\mathrm{W}^{1,p}(0,T)^{3}$. This solution has continuous mixed partial derivatives. From the uniqueness of the solution, we deduce the bijectivity of the function $X(t,\cdot)$, and its inverse is denoted by $Y(t,\cdot)$. Thanks to $\operatorname{div}b=0$, Liouville’s theorem yields that $X$ and $Y$ are both volume-preserving, $\mathrm{J}_{X}(t,y)\mathrm{J}_{Y}(t,X(t,y))=\operatorname{Id}_{3}$ and $\det\mathrm{J}_{X}(t,y)=\det\mathrm{J}_{Y}(t,x)=1$, with $\mathrm{J}_{X}$ and $\mathrm{J}_{Y}$ representing the Jacobian matrices of $X$ and $Y$ as usual. The inverse $Y$ of $X$ solves $\partial_{t}Y(t,x)=b^{(Y)}(t,Y(t,x))$ on $(0,T)\times\mathbb{R}^{3}$, where the right-hand side is given by $b^{(Y)}(t,y)\coloneqq-\mathrm{J}_{X}^{-1}(t,y)b(t,X(t,y))$, and $Y(0,x)=x$. Let us stress that $b^{(Y)}$ and $Y$ possess the same space and time regularity as $b$ and $X$. Moreover, the interval of existence of the solution is restricted by the condition $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in(0,T)$. In fact, we can write $\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq\|h(t)\|_{\mathbb{R}^{3}}+\|\mathbb{O}(t)-\mathbb{I}\|_{\mathbb{R}^{3\times 3}}|y|.$ As the rigid body moves with a continuous velocity, thanks to the estimates in Section 7, it is possible to find a sufficiently small time $T>0$ such that $\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq\frac{r}{2}$ is satisfied for large initial data. For small initial data, we will see that the exponential stability of the semigroup associated with the linearized problem at zero as revealed in Section 6 ensures that $\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq\frac{r}{2}$. Combining with the assumption $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$, we obtain $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in(0,\infty)$, so no collision occurs in finite time.
For $X\colon\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ solving the Cauchy problem (5.2) and $(t,y)\in(0,T)\times\mathbb{R}^{3}$, we define the transformed variables and stress tensor by
$v(t,y)$ $\coloneqq\mathrm{J}_{Y}(t,X(t,y))u^{\mathcal{F}}(t,X(t,y)),\kern 5.0ptQ(t,y)\coloneqq Q^{\mathcal{F}}(t,X(t,y)),\kern 5.0pt\ell(t)\coloneqq\mathbb{O}(t)^{\top}h^{\prime}(t),\kern 5.0pt\omega(t)\coloneqq\mathbb{O}(t)^{\top}\Omega(t),$
$\pi(t,y)$ $\coloneqq\pi^{\mathcal{F}}(t,X(t,y))\kern 5.0pt\text{and}\kern 5.0pt\Sigma(v(t,y),\pi(t,y),Q(t,y))\coloneqq\mathbb{O}(t)^{\top}\tilde{\Sigma}(\mathbb{O}(t)v(t,y),\pi(t,y),Q(t,y))\mathbb{O}(t).$
The metric contravariant $(g^{ij})$ and covariant tensors $(g_{ij})$ and the Christoffel symbol are given by
(5.3) $g^{ij}=\sum_{k=1}^{3}(\partial_{k}Y_{i})(\partial_{k}Y_{j}),\kern 5.0ptg_{ij}=\sum_{k=1}^{3}(\partial_{i}X_{k})(\partial_{j}X_{k})\kern 5.0pt\text{and}\kern 5.0pt\Gamma^{i}_{jk}=\frac{1}{2}\sum_{l=1}^{3}g^{il}(\partial_{k}g_{jl}+\partial_{j}g_{kl}-\partial_{l}g_{jk}).$
From [19, (3.13)] and [8, (3.13)], we recall the transformed Laplacians
(5.4) $(\mathcal{L}_{1}v)_{i}$ $=\sum_{j,k=1}^{3}\partial_{j}(g^{jk}\partial_{k}v_{i})+2\sum_{j,k,l=1}^{3}g^{kl}\Gamma_{jk}^{i}(\partial_{l}v_{j})+\sum_{j,k,l=1}^{3}\Bigl(\partial_{k}(g^{kl}\Gamma^{i}_{kl})+\sum_{m=1}^{3}g^{kl}\Gamma^{m}_{jl}\Gamma^{i}_{km}\Bigr)v_{j}\kern 5.0pt\text{and}\kern 5.0pt$
$(\mathcal{L}_{2}Q)_{ij}$ $=\sum\limits_{k,l=1}^{3}\left((\partial_{l}\partial_{k}Q_{ij})g^{kl}\right)+\sum\limits_{k=1}^{3}(\Delta Y_{k})(\partial_{k}Q_{ij}).$
Now, we compute the transformed time derivative, convective term and gradient of the pressure
$(\mathcal{M}v)_{i}=\sum\limits_{j=1}^{3}(\partial_{t}Y_{j})(\partial_{j}v_{i})+\sum\limits_{j,k=1}^{3}\bigl(\Gamma^{i}_{jk}(\partial_{t}Y_{k})+(\partial_{k}Y_{i})(\partial_{j}\partial_{t}X_{k})\bigr)v_{j},\kern 5.0pt(\mathcal{N}(v))_{i}=\sum\limits_{j=1}^{3}v_{j}(\partial_{j}v_{i})+\sum\limits_{j,k=1}^{3}\Gamma^{i}_{jk}v_{j}v_{k}$
and $(\mathcal{G}\pi)_{i}=\sum\limits_{j=1}^{3}g^{ij}(\partial_{j}\pi)$.
Before addressing the transformed terms associated to $\operatorname{div}\sigma$ and $\operatorname{div}\tau$, we settle some notation, where we omit the superscript F for simplicity. In that respect, we set $B_{\sigma}(Q)A$ so that it corresponds to $\operatorname{div}\sigma$, i. e., $B_{\sigma}(Q)Q=\operatorname{div}(Q\Delta Q-\Delta QQ)$. Thus, we define $B_{\sigma}(Q)A\coloneqq B_{\sigma,1}(Q)A+B_{\sigma,2}(Q)A$ by
(5.5) $B_{\sigma}(Q)A]_{i}\coloneqq\sum_{j,k=1}^{3}\Bigl(\bigl((\partial_{j}Q_{ik})(\Delta A_{kj})+Q_{ik}(\partial_{j}\Delta A_{kj})\bigr)-\bigl((\partial_{j}\Delta A_{ik})Q_{kj}+(\Delta A_{ik})(\partial_{j}Q_{kj})\bigr)\Bigr).$
The next term to be addressed is $\tau_{\mathrm{h}}$ from (3.1). We further split it into $\tau_{\mathrm{h}}\coloneqq\xi(\tau_{\mathrm{h},1}+\tau_{\mathrm{h},2}+\tau_{\mathrm{h},3})$, where $\tau_{\mathrm{h},1}\coloneqq-\nicefrac{{2}}{{3}}H_{\mathrm{h}}$, $\tau_{\mathrm{h},2}\coloneqq-\{Q,H_{\mathrm{h}}\}$ and $\tau_{\mathrm{h},3}\coloneqq 2(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(QH_{\mathrm{h}})$. We also introduce $B_{\tau_{\mathrm{h},i}}(Q)A$ such that $B_{\tau_{\mathrm{h},i}}(Q)Q=\operatorname{div}\tau_{\mathrm{h},i}$ for $i=1,2,3$. The first term is $B_{\tau_{\mathrm{h},1}}A]_{i}\coloneqq-\nicefrac{{2}}{{3}}\sum_{j=1}^{3}\partial_{j}(\Delta A_{ij}-A_{ij})$. Observing that $\{Q,H_{\mathrm{h}}\}=Q(\Delta Q)+(\Delta Q)Q-2Q^{2}$, we define
(5.6) $B_{\tau_{\mathrm{h},2}}(Q)A]_{i}\coloneqq-(B_{\sigma,1}(Q)A)_{i}+(B_{\sigma,2}(Q)A)_{i}+2\sum_{j,k=1}^{3}\bigl((\partial_{j}A_{ik})Q_{kj}+Q_{ik}(\partial_{j}A_{kj})\bigr).$
In the same way, calculating the term $\operatorname{tr}(QH_{\mathrm{h}})=\sum_{k,l=1}^{3}(Q_{kl}\Delta Q_{lk}-Q_{kl}Q_{lk})$, we introduce
i $\coloneqq 2\sum_{j,k,l=1}^{3}\Bigl((\partial_{j}Q_{ij})(Q_{kl}\Delta A_{lk}-Q_{kl}A_{lk})+(Q_{ij}+\nicefrac{{\delta_{ij}}}{{3}})\bigl((\partial_{j}Q_{kl})\Delta A_{lk}+Q_{kl}(\partial_{j}\Delta A_{lk})$
$\quad-(\partial_{j}Q_{kl})A_{lk}-Q_{kl}(\partial_{j}A_{lk})\bigr)\Bigr).$
We also write $\tau_{\mathrm{l}}=\xi\tau_{\mathrm{l},1}+\tau_{\mathrm{l},2}+\xi\tau_{\mathrm{l},3}$, with $\tau_{\mathrm{l},1}\coloneqq 2(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})(\operatorname{tr}(Q^{3})-\operatorname{tr}(Q^{2})^{2})$, $\tau_{\mathrm{l},2}\coloneqq-\nabla Q\odot\nabla Q$ and $\tau_{\mathrm{l},3}\coloneqq 2(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})H_{\mathrm{l}}$. Again, we define $B_{\tau_{\mathrm{l},i}}(Q)A$ so that $B_{\tau_{\mathrm{l},i}}(Q)Q=\operatorname{div}\tau_{\mathrm{l},i}$. The first term reads as
i $\coloneqq 2\sum_{j,k,l=1}^{3}\Biggl((\partial_{j}A_{ij})\biggl(\sum_{m=1}^{3}\Bigl(Q_{kl}Q_{lm}Q_{mk}-\sum_{n=1}^{3}Q_{km}Q_{mk}Q_{ln}Q_{nl}\Bigr)\biggr)+(Q_{ij}+\nicefrac{{\delta_{ij}}}{{3}})$
$\quad\cdot\biggl(\sum_{m=1}^{3}\Bigl((\partial_{j}A_{kl})Q_{lm}Q_{mk}+Q_{kl}(\partial_{j}A_{lm})Q_{mk}+Q_{kl}Q_{lm}(\partial_{j}A_{mk})-\sum_{n=1}^{3}\bigl((\partial_{j}A_{km})Q_{mk}Q_{ln}Q_{nl}$
$\qquad+Q_{km}(\partial_{j}A_{mk})Q_{ln}Q_{nl}+Q_{km}Q_{mk}(\partial_{j}A_{ln})Q_{nl}+Q_{km}Q_{mk}Q_{ln}(\partial_{j}A_{nl})\bigr)\Bigr)\biggr)\Biggr).$
To capture $\nabla Q\odot\nabla Q$, we set $B_{\tau_{\mathrm{l},2}}(Q)A]_{i}\coloneqq-\sum_{j,k=1}^{2}\bigl((\partial_{i}Q_{jk})\Delta A_{kj}+\sum_{l=1}^{3}(\partial_{l}Q_{kj})(\partial_{l}\partial_{i}A_{jk})\bigr)$. The term $B_{\tau_{\mathrm{l},3}}(Q)A$ associated to $\tau_{\mathrm{l},3}$ is of a similar form as $B_{\tau_{\mathrm{l},1}}(Q)A$.
Next, we calculate the respective transformed terms. The transformed version of $B_{\sigma}$ from (5.5) is
(5.7) $\mathcal{B}_{\sigma}(Q)A\coloneqq\mathcal{B}_{\sigma,1}(Q)A+\mathcal{B}_{\sigma,2}(Q)A\coloneqq\mathcal{B}_{\sigma,11}(Q)A+\mathcal{B}_{\sigma,12}(Q)A+\mathcal{B}_{\sigma,2}(Q)A,\kern 5.0pt\text{where}\kern 5.0pt$
$(\mathcal{B}_{\sigma,11}(Q)A)_{i}$ $\coloneqq\sum_{j,k=1}^{3}(\mathcal{L}_{2}A)_{kj}\sum_{l=1}^{3}(\partial_{l}Q_{ik})(\partial_{j}Y_{l}),\kern 5.0pt(\mathcal{B}_{\sigma,12}(Q)A)_{i}\coloneqq\sum_{j,k=1}^{3}Q_{ik}\Biggl(\sum_{l=1}^{3}\biggl(\sum_{m=1}^{3}\Bigl(\sum_{n=1}^{3}(\partial_{n}\partial_{m}\partial_{l}A_{kj})$
$\quad\cdot(\partial_{j}Y_{n})g^{lm}\Bigr)+(\partial_{m}\partial_{l}A_{kj})(\partial_{j}g^{lm})+(\partial_{m}\partial_{l}A_{kj})(\partial_{j}Y_{m})\Delta Y_{l}\biggr)+(\partial_{l}A_{kj})(\partial_{j}\Delta Y_{l})\Biggr).$
By (5.5), the term $\mathcal{B}_{\sigma,2}(Q)A$ takes a similar shape. For the transformed versions of $B_{\tau_{\mathrm{h},i}}$, we first obtain
(5.8) $(\mathcal{B}_{\tau_{\mathrm{h},1}}A)_{i}$ $=-\frac{2}{3}\sum_{j,k=1}^{3}\biggl(\sum_{l=1}^{3}\Bigl(\sum_{m=1}^{3}\bigl((\partial_{m}\partial_{l}\partial_{k}A_{ij})(\partial_{j}Y_{m})g^{kl}\bigr)+(\partial_{l}\partial_{k}A_{ij})(\partial_{j}g^{kl})$
$\qquad+(\partial_{l}\partial_{k}A_{ij})(\partial_{j}Y_{l})\Delta Y_{k}\Bigr)+(\partial_{k}A_{ij})(\partial_{j}\Delta Y_{k})+(\partial_{j}Y_{k})(\partial_{k}A_{ij})\biggr).$
From (5.6), it follows that
$(\mathcal{B}_{\tau_{\mathrm{h},2}}(Q)A)_{i}=-(\mathcal{B}_{\sigma,1}(Q)A)_{i}+(\mathcal{B}_{\sigma,2}(Q)A)_{i}+2\sum_{j,k,l=1}^{3}\bigl((\partial_{j}Y_{l})(\partial_{l}A_{ik})Q_{kj}+Q_{ik}(\partial_{j}Y_{l})(\partial_{l}A_{kj})\bigr).$
The $i$-th component of the last term $\mathcal{B}_{\tau_{\mathrm{h},3}}(Q)A$ in this context is given by
$(\mathcal{B}_{\tau_{\mathrm{h},3}}(Q)A)_{i}$ $=2\sum_{j,k,l,m=1}^{3}\biggl((\partial_{j}Y_{m})(\partial_{m}Q_{ij})(Q_{kl}(\mathcal{L}_{2}A)_{lk}-Q_{kl}A_{lk})+(Q_{ij}+\nicefrac{{\delta_{ij}}}{{3}})\Bigl((\partial_{j}Y_{m})(\partial_{m}Q_{kl})(\mathcal{L}_{2}A)_{lk}$
$\quad+(\mathcal{B}_{\sigma,12}(Q)A)_{j}-(\partial_{j}Y_{m})(\partial_{m}Q_{kl})A_{lk}-Q_{kl}(\partial_{j}Y_{m})(\partial_{m}A_{lk})\Bigr)\biggr).$
Similarly, we obtain the transformed terms corresponding to $B_{\tau_{\mathrm{l}}}$. For brevity, we omit the details. Let us summarize the transformed terms given above in a more compact form as
$\mathcal{B}_{\tau_{\mathrm{h}}}(Q)A$ $\coloneqq\xi(\mathcal{B}_{\tau_{\mathrm{h},1}}A+\mathcal{B}_{\tau_{\mathrm{h},2}}(Q)A+\mathcal{B}_{\tau_{\mathrm{h},3}}(Q)A),\kern 5.0pt\mathcal{B}_{\tau_{\mathrm{l}}}(Q)A\coloneqq\xi\mathcal{B}_{\tau_{\mathrm{l},1}}(Q)A+\mathcal{B}_{\tau_{\mathrm{l},2}}(Q)A+\xi\mathcal{B}_{\tau_{\mathrm{l},3}}(Q)A,$
$\mathcal{B}_{\tau}(Q)A$ $\coloneqq\mathcal{B}_{\tau_{\mathrm{h}}}(Q)A+\mathcal{B}_{\tau_{\mathrm{l}}}(Q)A\kern 5.0pt\text{and}\kern 5.0pt\mathcal{B}_{\tau}(Q)\coloneqq\mathcal{B}_{\tau}(Q)Q\kern 5.0pt\text{as well as}\kern 5.0pt\mathcal{B}_{\sigma}(Q)\coloneqq\mathcal{B}_{\sigma}(Q)Q.$
In the equation for $Q$, we first observe that $\partial_{t}Q^{\mathcal{F}}=\partial_{t}Q+(\partial_{t}Y\cdot\nabla)Q$. For the convective term, we obtain $(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}=\sum_{k=1}^{3}\bigl(\sum_{l=1}^{3}(\partial_{l}X_{k})v_{l}\bigr)\bigl(\sum_{m=1}^{3}(\partial_{m}Q_{ij})(\partial_{k}Y_{m})\bigr)=\sum_{l,m=1}^{3}\delta_{lm}v_{l}(\partial_{m}Q_{ij})=(v\cdot\nabla)Q$.
Concerning the transformation of the gradient of $u^{\mathcal{F}}$, we calculate
(5.9) $\frac{\partial u_{i}^{\mathcal{F}}}{\partial x_{j}}=\sum_{k=1}^{3}\Bigl(\sum_{l=1}^{3}(\partial_{l}\partial_{k}X_{i})(\partial_{j}Y_{l})v_{k}+(\partial_{k}X_{i})\sum_{l=1}^{3}(\partial_{l}v_{k})(\partial_{j}Y_{l})\Bigr)\eqqcolon\mathcal{D}(v)_{ij}\eqqcolon\mathcal{D}_{ij}.$
Consequently, $S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})$ from (3.1) transforms as
(5.10) $(\mathcal{S}(v,Q))_{ij}$ $=-\frac{1}{2}\sum_{k=1}^{3}\bigl(Q_{ik}(\mathcal{D}_{kj}-\mathcal{D}_{jk})-(\mathcal{D}_{ik}-\mathcal{D}_{ki})Q_{kj}\bigr)+\xi\biggl(\frac{1}{3}(\mathcal{D}_{ij}+\mathcal{D}_{ji})$
$\qquad+\frac{1}{2}\sum_{k=1}^{3}Q_{ik}(\mathcal{D}_{kj}+\mathcal{D}_{jk})+(\mathcal{D}_{ik}+\mathcal{D}_{ki})Q_{kj}-2\Bigl(\sum_{k,l=1}^{3}Q_{kl}\mathcal{D}_{lk}\Bigr)\Bigl[Q_{ij}+\frac{1}{3}\delta_{ij}\Bigr]\biggr).$
The term $H(Q)$ from (2.3) transforms to $\mathcal{H}(Q)=\mathcal{L}_{2}Q-Q+(Q^{2}-\operatorname{tr}(Q^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}})-\operatorname{tr}(Q^{2})Q$. Then for the unit outward normal $n=\mathbb{O}^{\top}(t)\nu(t)$ to $\partial\mathcal{F}_{0}$, which is thus directed inwards to $\partial\mathcal{S}_{0}$, the complete transformed system associated to (2.5)–(2.9) on the fixed domain $(0,T)\times\mathcal{F}_{0}$ takes the shape
(5.11) $\left\{\begin{aligned} \partial_{t}v+(\mathcal{M}-\mathcal{L}_{1})v+\mathcal{N}(v)+\mathcal{G}\pi&=\mathcal{B}_{\tau}(Q)+\mathcal{B}_{\sigma}(Q),\kern 5.0pt\operatorname{div}v=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}Q}+(\partial_{t}Y\cdot\nabla)Q+(v\cdot\nabla)Q-\mathcal{S}(v,Q)&=\mathcal{H}(Q),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}\ell^{\prime}+m_{\mathcal{S}}(\omega\times\ell)&=-\int\limits_{\partial\mathcal{S}_{0}}\Sigma(v,\pi,Q)n\,\mathrm{d}\Gamma,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ J_{0}\omega^{\prime}-\omega\times(J_{0}\omega)&=-\int\limits_{\partial\mathcal{S}_{0}}y\times\Sigma(v,\pi,Q)n\,\mathrm{d}\Gamma,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T).\end{aligned}\right$.
The boundary and initial conditions are
(5.12) $v=0,\kern 5.0pt\partial_{n}Q=0\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0ptv=\ell+\omega\times y,\kern 5.0pt\partial_{n}Q=0\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{S}_{0}\kern 5.0pt\text{and}\kern 5.0pt$
(5.13) $v(0)=v_{0},\kern 5.0ptQ(0)=Q_{0},\kern 5.0pt\ell(0)=\ell_{0}\kern 5.0pt\text{and}\kern 5.0pt\omega(0)=\omega_{0}.$
Thanks to the above coordinate transform, we may rewrite the two main results, Theorem 4.1 and Theorem 4.2, on the fixed domain $(0,T)\times\mathcal{F}_{0}$, which we also refer to as the reference configuration.
Theorem 5.1.
Consider $p>4$, a bounded $\mathrm{C}^{3}$-domain $\mathcal{O}\subset\mathbb{R}^{3}$, the initial rigid body and fluid domains $\mathcal{S}_{0}$ and $\mathcal{F}_{0}$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, with $\mathrm{X}_{\gamma}$ as in (2.10). If there is $r>0$ with $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$, then there exist $T^{\prime}\in(0,T]$ and $X\in\mathrm{C}^{1}([0,T^{\prime}];\mathrm{C}^{2}(\mathbb{R}^{3})^{3})$ so that $X(\tau,\cdot)\colon\mathcal{F}(\tau)\to\mathcal{F}_{0}$ are $\mathrm{C}^{2}$-diffeomorphisms for all $\tau\in[0,T^{\prime}]$, and the interaction problem (5.11)–(5.13) has a unique solution
$v$ $\in\mathrm{W}^{1,p}\bigl(0,T^{\prime};\mathrm{L}^{2}(\mathcal{F}_{0})^{3}\bigr)\cap\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\bigr),\kern 5.0ptQ\in\mathrm{W}^{1,p}\bigl(0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr)\cap\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr),$
$\ell$ $\in\mathrm{W}^{1,p}(0,T^{\prime})^{3},\kern 5.0pt\omega\in\mathrm{W}^{1,p}(0,T^{\prime})^{3}\kern 5.0pt\text{and}\kern 5.0pt\pi\in\mathrm{L}^{p}\bigl(0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})\bigr).$
Theorem 4.2 can be reformulated in the following manner on the reference configuration.
Theorem 5.2.
Let $p>4$, let $\mathcal{O}\subset\mathbb{R}^{3}$ be a bounded domain of class $\mathrm{C}^{3}$, and consider the domains of the rigid body and the fluid at time zero $\mathcal{S}_{0}$ and $\mathcal{F}_{0}$. Besides, assume that $\mathrm{dist}(\mathcal{S}_{0},\partial\mathcal{O})>r$ for some $r>0$. Then there exists $\eta_{0}>0$ such that for all $\eta\in(0,\eta_{0})$, there are $\delta_{0}>0$, $C>0$ so that for every $\delta\in(0,\delta_{0})$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\overline{B}_{\mathrm{X}_{\gamma}}(0,\delta)$, it holds that $X\in\mathrm{C}^{1}([0,\infty);\mathrm{C}^{2}(\mathbb{R}^{3})^{3})$, with $X(\tau,\cdot)\colon\mathcal{F}(\tau)\to\mathcal{F}_{0}$ being $\mathrm{C}^{2}$-diffeomorphisms for all $\tau\in[0,\infty)$, and (5.11)–(5.13) has a unique solution $(v,Q,\ell,\omega,\pi)$ with
$\|\mathrm{e}^{\eta(\cdot)}v\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{2}(\mathcal{F}_{0}))}+\|\mathrm{e}^{\eta(\cdot)}\partial_{t}v\|_{\mathrm{L}^{p}(0,\infty;\mathrm{L}^{2}(\mathcal{F}_{0}))}+\|\mathrm{e}^{\eta(\cdot)}v\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{B}_{2p}^{2-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0}))}$
$+\|\mathrm{e}^{\eta(\cdot)}Q\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{3}(\mathcal{F}_{0}))}+\|\mathrm{e}^{\eta(\cdot)}\partial_{t}Q\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{1}(\mathcal{F}_{0}))}+\|\mathrm{e}^{\eta(\cdot)}Q\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0}))}$
$+\|\mathrm{e}^{\eta(\cdot)}\ell\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\omega\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\pi\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0}))}\leq C\delta.$
6. Maximal regularity of the linearized problem and exponential stability
This section is central for the article. Very recently, the sectoriality of the linearized general $Q$-tensor problem has been shown in [22]. We take advantage of this property to derive the maximal $\mathrm{L}^{p}$-regularity of the suitably linearized interaction problem by introducing a so-called fluid-structure operator tailored to our setting. In a second step, we prove the exponential stability of the corresponding semigroup.
Let us first introduce some function spaces. For $J=(0,T)$, $0<T\leq\infty$, the data space is defined by
(6.1) $\mathbb{F}_{T}=\mathbb{F}_{T}^{v}\times\mathbb{F}_{T}^{Q}\times\mathbb{F}_{T}^{\ell}\times\mathbb{F}_{T}^{\omega}=\mathrm{L}^{p}\bigl(J;\mathrm{L}^{2}(\mathcal{F}_{0})^{3}\bigr)\times\mathrm{L}^{p}\bigl(J;\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr)\times\mathrm{L}^{p}(J)^{3}\times\mathrm{L}^{p}(J)^{3}.$
Next, with $z=(v,Q,\ell,\omega)$, we introduce $\mathrm{X}_{0}\coloneqq\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}$ and
$\mathrm{X}_{1}\coloneqq\{z\in\mathrm{Y}_{1}:v=0,\kern 5.0pt\partial_{n}Q=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{O},\kern 5.0pt\text{and}\kern 5.0ptv=\ell+\omega\times y,\kern 5.0pt\partial_{n}Q=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{0}\}$
for $\mathrm{Y}_{1}\coloneqq\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\cap\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}$. As in [8, Lemma 5.3], $\mathrm{X}_{\gamma}$ from (2.10) can be deduced from $\mathrm{X}_{0}$ and $\mathrm{X}_{1}$ by real interpolation via $\mathrm{X}_{\gamma}=(\mathrm{X}_{0},\mathrm{X}_{1})_{1-\nicefrac{{1}}{{p}},p}$. We then define the solution space $\mathbb{E}_{T}$ by
(6.2) $\mathbb{E}_{T}\coloneqq\mathrm{W}^{1,p}(J;\mathrm{X}_{0})\cap\mathrm{L}^{p}(J;\mathrm{X}_{1}).$
By $\prescript{}{0}{\mathbb{E}_{T}}$, we denote the associated space with homogeneous initial values. For simplicity, we also write $\mathbb{E}_{T}^{Q}$ for the $Q$-component of $\mathbb{E}_{T}$. In order to account for the pressure, we further introduce
(6.3) $\tilde{\mathbb{E}}_{T}\coloneqq\mathbb{E}_{T}\times\mathbb{E}_{T}^{\pi},\kern 5.0pt\text{where}\kern 5.0pt\mathbb{E}_{T}^{\pi}\coloneqq\mathrm{L}^{p}\bigl(J;\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})\bigr),$
and $\|\cdot\|_{\mathbb{E}_{T}^{\pi}}$ represents the norm associated to $\mathbb{E}_{T}^{\pi}$. With regard to the stress tensor part in the surface integral, we include precisely its linear components to establish estimates in the proof of Theorem 4.2 in Section 8. In addition to the usual Cauchy stress tensor, we thus incorporate $\nicefrac{{2\xi}}{{3}}(\Delta Q-Q)$, i. e.,
(6.4) $\mathrm{T}(v,\pi,Q)\coloneqq\nabla v+(\nabla v)^{\top}-\pi\operatorname{Id}_{3}-\frac{2\xi}{3}(\Delta Q-Q).$
We need to include the $Q$-terms in order to obtain the exponential stability for the linearization at zero, and to establish the estimates of the surface integrals involving the stress tensor.
As the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity already requires the associated operator to be invertible, we introduce a shift, i. e., instead of considering the original linearized problem, we add an artificial term of the form $\mu\operatorname{Id}$ on the left-hand side of the linearized problem, see (6.5) below. It does not pose any problems for the local well-posedness, but it is only critical for the global well-posedness. Thus, for $\mu\geq 0$ as well as $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, $(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{T}$ and $(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, recalling $S_{\xi}$ and $\tilde{S}_{\xi}$ from (3.2), we consider the linearization
(6.5) $\left\{\begin{aligned} \partial_{t}{v}+(\mu-\Delta){v}+\nabla\pi-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\kern 5.0pt\operatorname{div}{v}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(\widehat{Q})\nabla v+(\mu-(\Delta-\operatorname{Id}_{3})){Q}&=f_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\ell})^{\prime}+\mu\ell+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=f_{3},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ J_{0}({\omega})^{\prime}+\mu\omega+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=f_{4},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ {v}={\ell}+{\omega}\times y,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{S}_{0},\kern 5.0pt{v}=0,&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0pt\partial_{n}Q=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\kern 5.0ptQ(0)=Q_{0},\kern 5.0pt\ell(0)&=\ell_{0}\kern 5.0pt\text{and}\kern 5.0pt\omega(0)=\omega_{0}.\end{aligned}\right$.
Below, we assert the main result of this section on the maximal $\mathrm{L}^{p}$-regularity of (6.5), up to a shift by $\mu>0$ for general $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$ and without shift in the case $\widehat{Q}=0$. Even though we only require this result for $p>4$, we provide the more general case $p\in(1,\infty)\setminus\{\nicefrac{{4}}{{3}}\}$ here. We do not need a more restrictive condition on $p$, because we consider a linearized problem and take into account $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$.
Theorem 6.1.
Let $p\in(1,\infty)\setminus\{\nicefrac{{4}}{{3}}\}$, $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, $T\in(0,\infty]$, $(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{T}$ as well as $(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$.
  1. (a)
    There is $\mu_{0}\geq 0$ such that for all $\mu>\mu_{0}$, (6.5) admits a unique solution $(v,Q,\ell,\omega,\pi)\in\tilde{\mathbb{E}}_{T}$, and there exists $C_{\mathrm{MR}}>0$ such that $\|(v,Q,\ell,\omega,\pi)\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}\bigl(\|(f_{1},f_{2},f_{3},f_{4})\|_{\mathbb{F}_{T}}+\|(v_{0},Q_{0},\ell_{0},\omega_{0})\|_{\mathrm{X}_{\gamma}}\bigr)$. The constant $C_{\mathrm{MR}}>0$ can be chosen independent of $T$ in the case of homogeneous initial values.
  2. (b)
    If $\widehat{Q}=0$, then the assertion from (a) is also valid without shift, that means, there exists a unique solution $(v,Q,\ell,\omega,\pi)\in\tilde{\mathbb{E}}_{T}$ to (6.5) for $\mu=0$. In particular, there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0})$, if $\mathrm{e}^{\eta t}(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{\infty}$, then there is a unique solution $\mathrm{e}^{\eta t}(v,Q,\ell,\omega,\pi)\in\tilde{\mathbb{E}}_{\infty}$ to (6.5) with $\mu=0$. Besides, there exists a constant $C_{\mathrm{MR}}>0$ such that
    $\quad\|\mathrm{e}^{\eta(\cdot)}v\|_{\mathbb{E}_{\infty}^{v}}+\|\mathrm{e}^{\eta(\cdot)}\pi\|_{\mathbb{E}_{\infty}^{\pi}}+\|\mathrm{e}^{\eta(\cdot)}Q\|_{\mathbb{E}_{\infty}^{Q}}+\|\mathrm{e}^{\eta(\cdot)}\ell\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\omega\|_{\mathrm{W}^{1,p}(0,\infty)}$
    $\leq C_{\mathrm{MR}}\bigl(\|(v_{0},Q_{0},\ell_{0},\omega_{0})\|_{\mathrm{X}_{\gamma}}+\|\mathrm{e}^{\eta(\cdot)}(f_{1},f_{2},f_{3},f_{4})\|_{\mathbb{F}_{\infty}}\bigr).$
A remark on extensions of the Hilbert space case is in order.
Remark 6.2.
From [3, Proposition 7.17], the maximal $\mathrm{L}^{p}$-regularity of the linearized liquid crystal problem on $\mathrm{L}^{q}\times\mathrm{W}^{1,q}$ follows for $q$ close to $2$. Thus, we expect that an extension of Theorem 6.1 to $q$ close to $2$ is possible. The case of general $q\in(1,\infty)$, however, remains an open problem.
The rest of this section is dedicated to the proof of Theorem 6.1, and the main idea is to reformulate (6.5) in operator form by invoking an adapted fluid-structure operator. At first, we consider the $Q$-tensor part of the linearized problem (6.5) with homogeneous boundary conditions. For simplicity, we omit the shift, while the interface condition is handled by a lifting procedure later on. We denote by $u=(v,Q)$ the principle variable. For suitable $f_{1}$, $f_{2}$, $v_{0}$, $Q_{0}$, and for $S_{\xi}$ and $\tilde{S}_{\xi}$ from (3.2), the resulting problem is
(6.6) $\left\{\begin{aligned} \partial_{t}{v}-\Delta{v}+\nabla\pi-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\kern 5.0pt\operatorname{div}{v}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(\widehat{Q})\nabla v-(\Delta-\operatorname{Id}_{3}){Q}&=f_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {v}=0,\kern 5.0pt\partial_{n}Q&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\kern 5.0ptQ(0)&=Q_{0},&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathcal{F}_{0}.\end{aligned}\right$.
Following [22, Section 5], we introduce the operator corresponding to (6.6). First, we recall the Helmholtz projection $\mathbb{P}\colon\mathrm{L}^{2}(\mathcal{F}_{0})^{3}\to\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ on $\mathrm{L}^{2}$ and invoke the Stokes operator $A_{2}$ on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ defined by
(6.7) $A_{2}v\coloneqq\mathbb{P}\Delta v,\kern 5.0pt\text{for}\kern 5.0ptv\in\mathrm{D}(A_{2})\coloneqq\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\cap\mathrm{H}_{0}^{1}(\mathcal{F}_{0})^{3}\cap\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0}).$
Besides, the modified Neumann-Laplacian on $\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})$ is defined by
$\Delta_{\mathrm{N}}^{1}Q\coloneqq(\Delta-\operatorname{Id}_{3})Q,\kern 5.0pt\text{for}\kern 5.0ptQ\in\mathrm{D}(\Delta_{\mathrm{N}}^{1})\coloneqq\{Q\in\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3}):\partial_{n}Q=0\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{F}_{0}\}.$
We introduce $\mathrm{X}_{0}^{\mathrm{lc}}\coloneqq\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})$ as well as $\mathrm{X}_{1}^{\mathrm{lc}}\coloneqq\mathrm{D}(A_{2})\times\mathrm{D}(\Delta_{\mathrm{N}}^{1})$ for the lemma below. Given $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, we define the operator matrix $A_{\xi}(\widehat{Q})\colon\mathrm{X}_{1}^{\mathrm{lc}}\to\mathrm{X}_{0}^{\mathrm{lc}}$ by
(6.8) $A_{\xi}(\widehat{Q})\coloneqq\begin{pmatrix}-A_{2}&-\mathbb{P}\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})\\ \tilde{S}_{\xi}(\widehat{Q})\nabla&-\Delta_{\mathrm{N}}^{1}\end{pmatrix}$.
We collect auxiliary results for our analysis in the lemma below. Part (a) will be an ingredient in the proof of the sectoriality of the operator associated with the linearized interaction problem, while part (b) will be used for the spectral analysis.
Lemma 6.3 ([22, Section 5]).
Let $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$ and $u=(v,Q)\in\mathrm{X}_{1}^{\mathrm{lc}}$. Then
  1. (a)
[…]
  2. (b)
    $\langle A_{\xi}(\widehat{Q})u,u\rangle_{\mathrm{X}_{0}^{\mathrm{lc}}}=\|\nabla v\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\|Q\|_{\mathrm{H}^{2}(\mathcal{F}_{0})}^{2}+\|\nabla Q\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\mathrm{i}\operatorname{Im}(2\langle\nabla v,S_{\xi}(\widehat{Q})(-\Delta+\operatorname{Id}_{3})Q\rangle_{\mathrm{L}^{2}(\mathcal{F}_{0})}$.
6.1. Lifting operators and added mass
In this subsection, we address the stress tensor part in the surface integrals. It is sufficient to concentrate on the linearized interaction problem of the fluid part at first and to handle then the $Q$-term in (6.4) by means of perturbation theory. Hence, to shorten notation, we introduce the Cauchy stress tensor
(6.9) $\tilde{\mathrm{T}}(v,\pi)\coloneqq\nabla v+(\nabla v)^{\top}-\pi\operatorname{Id}_{3}.$
A key concept is the added mass operator, see e. g. [17] or [19, Section 4.2] for further background. The most important aspects about this approach are the existence of lifting operators for the inhomogeneous interface condition and a representation formula of the pressure appearing in the surface integrals. This will be of central relevance for the reformulation of (6.5) in operator form. Let us consider
(6.10) $\left\{\begin{aligned} -\Delta w+\nabla\psi&=0,\kern 5.0pt\operatorname{div}w=0,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathcal{F}_{0},\\ w&=0,\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{O},\kern 5.0ptw=\ell+\omega\times y,&&\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{S}_{0},\\ \int_{\mathcal{F}_{0}}\psi\,\mathrm{d}y&=0.\end{aligned}\right$.
The following result on the solvability of (6.10) is classical. We refer for instance to [27, Lemma 3.6]. It provides an elementary representation of the solution to the Stokes problem with inhomogeneous boundary conditions at the interface, capturing the motion of the rigid body.
Lemma 6.4.
Consider $(\ell,\omega)\in\mathbb{C}^{3}\times\mathbb{C}^{3}$ and the canonical basis $\{e_{i}\}$ in $\mathbb{C}^{3}$. Then the solution $(w,\psi)$ to (6.10) admits the representation $w=\sum_{i=1}^{3}\ell_{i}W_{i}+\sum_{i=4}^{6}\omega_{i-3}W_{i}$ and $\psi=\sum_{i=1}^{3}\ell_{i}\Psi_{i}+\sum_{i=4}^{6}\omega_{i-3}\Psi_{i}$, where $(W_{i},\Psi_{i})$, $i=1,\dots,6$, in turn satisfy $\int_{\mathcal{F}_{0}}\Psi_{i}\,\mathrm{d}y=0$ and
$\left\{\begin{aligned} -\Delta W_{i}+\nabla\Psi&=0,\kern 5.0pt\operatorname{div}W_{i}=0,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathcal{F}_{0},\\ W_{i}&=0,&&\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{O},\\ W_{i}=e_{i},\kern 5.0pt\text{for}\kern 5.0pti=1,2,3\kern 5.0pt\text{and}\kern 5.0ptW_{i}&=e_{i-3}\times y,\kern 5.0pt\text{for}\kern 5.0pti=4,5,6,&&\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{S}_{0},\end{aligned}\right$.
$\mathbb{B}_{i,j}\coloneqq\int_{\mathcal{F}_{0}}DW_{i}:DW_{j}\,\mathrm{d}y$ is invertible, where $DW_{i}$ represents the gradient of $W_{i}$, and for $\tilde{\mathrm{T}}$ from (6.9), we have $\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}\tilde{\mathrm{T}}(w,\psi)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times\tilde{\mathrm{T}}(w,\psi)n\,\mathrm{d}\Gamma\end{pmatrix}=\mathbb{B}\binom{\ell}{\omega}$.
Subsection 6.1 allows us to define lifting operators $D_{\mathrm{fl}}\in\mathcal{L}(\mathbb{C}^{6},\mathrm{H}^{2}(\mathcal{F}_{0}))$ and $D_{\mathrm{pr}}\in\mathcal{L}(\mathbb{C}^{6},\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0}))$. In fact, given $(\ell,\omega)\in\mathbb{C}^{3}\times\mathbb{C}^{3}$, for the solution $(w,\psi)$ to (6.10) emerging from Subsection 6.1, we set
(6.11) $D_{\mathrm{fl}}(\ell,\omega)\coloneqq w\kern 5.0pt\text{and}\kern 5.0ptD_{\mathrm{pr}}(\ell,\omega)\coloneqq\psi.$
We also invoke the operator $N\in\mathcal{L}\bigl(\mathrm{H}^{\nicefrac{{1}}{{2}}}(\partial\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\partial\mathcal{F}_{0}),\mathrm{H}^{2}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})\bigr)$, with $Nh\coloneqq\varphi$, for $\varphi$ solving
(6.12) $\Delta\varphi=0,\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{0},\kern 5.0pt\partial_{n}\varphi=h,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{F}_{0}.$
For the Neumann operator $N$ and $h\in\mathrm{H}^{\nicefrac{{1}}{{2}}}(\partial\mathcal{S}_{0})\cap\mathrm{L}_{0}^{2}(\partial\mathcal{S}_{0})$, we define the operator $N_{\mathcal{S}_{0}}$ by
(6.13) $N_{\mathcal{S}_{0}}h\coloneqq N(\chi_{\partial\mathcal{S}_{0}}h).$
The above lifting operators allow us to rewrite the resolvent problem linked to $v$ in terms of $\mathbb{P}v$ and $(\operatorname{Id}_{3}-\mathbb{P})v$ for the Helmholtz projection $\mathbb{P}$ on $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. Besides, we obtain a formula for the pressure in terms of $\mathbb{P}v$ as well as the rigid body velocities. This will be important for handling the pressure term appearing in the surface integrals of the stress tensor, and it paves the way for a reformulation in operator form. This splitting idea for fluid-structure interaction problems was e. g. used by Raymond [36]. For a proof of the next lemma, we refer to [27, Proposition 3.7].
Lemma 6.5.
Assume $(f_{1},\ell,\omega)\in\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathbb{C}^{3}\times\mathbb{C}^{3}$, and recall the Stokes operator on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ from (6.7), the Dirichlet lifting operator $D_{\mathrm{fl}}$ from (6.11), the lifting operator $N$ associated to the Neumann problem (6.12) as well as $N_{\mathcal{S}_{0}}$ from (6.13). Then $(v,\pi)\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ satisfies
$\left\{\begin{aligned} \lambda v-\Delta v+\nabla\pi&=f_{1},\kern 5.0pt\operatorname{div}v=0,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathcal{F}_{0},\\ v=0,\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{O},\kern 5.0ptv&=\ell+\omega\times y,&&\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{S}_{0},\end{aligned}\quad\text{if and only if}\right$.
(6.14) $\left\{\begin{aligned} \lambda\mathbb{P}v-A_{2}\mathbb{P}v+A_{2}\mathbb{P}D_{\mathrm{fl}}(\ell,\omega)&=\mathbb{P}f_{1},\\ (\operatorname{Id}_{3}-\mathbb{P})v&=(\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega),\\ \pi&=N(\Delta\mathbb{P}v\cdot n)-\lambda N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n).\end{aligned}\right$.
For the reformulation of (6.5) in operator form, we invoke the corresponding resolvent problem. Indeed, for $\lambda\in\mathbb{C}$ and $(f_{1},f_{2},f_{3},f_{4})\in\mathrm{L}^{2}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{C}^{3}\times\mathbb{C}^{3}$, it reads as
(6.15) $\left\{\begin{aligned} \lambda v-\Delta v+\nabla\pi-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\kern 5.0pt\operatorname{div}v=0,&&\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{0},\\ \lambda Q+\tilde{S}_{\xi}(\widehat{Q})\nabla v-(\Delta-\operatorname{Id})Q&=f_{2},&&\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{0},\\ \lambda m_{\mathcal{S}}\ell+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=f_{3},\\ \lambda J_{0}\omega+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=f_{4},\\ v=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{O},\kern 5.0ptv&=\ell+\omega\times y,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{0},\kern 5.0pt\partial_{n}Q=0,&&\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{F}_{0}.\end{aligned}\right$.
Thanks to (6.14), we rewrite (6.15)3 as $\lambda K\binom{\ell}{\omega}=C_{1}\mathbb{P}v+C_{2}Q+C_{3}\binom{\ell}{\omega}+\binom{f_{3}}{f_{4}}$, where
(6.16) $K$ $=\begin{pmatrix}m_{\mathcal{S}}\operatorname{Id}_{3}&0\\ 0&J_{0}\end{pmatrix}+M,\kern 5.0pt\text{with}\kern 5.0ptM\binom{\ell}{\omega}=\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n)n\,\mathrm{d}\Gamma\end{pmatrix}$,
$C_{1}\mathbb{P}v$ $=-2\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}\mathbb{D}(\mathbb{P}v)n\,\mathrm{d}\Gamma+\int_{\partial\mathcal{S}_{0}}N(\Delta\mathbb{P}v\cdot n)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times\mathbb{D}(\mathbb{P}v)n\,\mathrm{d}\Gamma+\int_{\partial\mathcal{S}_{0}}y\times N(\Delta\mathbb{P}v\cdot n)n\,\mathrm{d}\Gamma\end{pmatrix}$,
$C_{2}Q$ $=\frac{2\xi}{3}\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}(\Delta Q-Q)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times(\Delta Q-Q)n\,\mathrm{d}\Gamma\end{pmatrix},\kern 5.0ptC_{3}\binom{\ell}{\omega}=-2\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}\mathbb{D}((\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega))n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times\mathbb{D}((\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega))n\,\mathrm{d}\Gamma\end{pmatrix}$.
Let us note that $M$ above is generally referred to as the added mass operator. From, [17, Lemma 4.6], see also [19, Lemma 4.3], we recall that the matrix $K$ from (6.16) is invertible. The purpose of the added mass operator is that allows for a reformulation of the surface integrals without the pressure. More precisely, thanks to the invertibility of $K$, it is possible to reformulate the equations for the rigid body velocities including the surface integrals as a problem of $\mathbb{P}v$, $Q$ and a lifted problem involving the rigid body velocities $\ell$ and $\omega$ themselves.
6.2. The fluid-structure operator
Thanks to the operators introduced in the previous subsection, we are now in the position to define the fluid-structure operator corresponding to the linearized $Q$-tensor interaction problem (6.5). For $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, the operator $A_{\mathrm{fs}}(\widehat{Q})\colon\mathrm{D}(A_{\mathrm{fs}})\subset\mathrm{X}\to\mathrm{X}$ is defined on $\mathrm{X}\coloneqq\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{C}^{3}\times\mathbb{C}^{3}$, and with domain $\mathrm{D}(A_{\mathrm{fs}})\coloneqq\{(\mathbb{P}v,Q,\ell,\omega)\in\mathrm{X}:\mathbb{P}v-\mathbb{P}D_{\mathrm{fl}}(\ell,\omega)\in\mathrm{D}(A_{2})\}$ it takes the shape
[…]
The next proposition reveals that the resolvent problem (6.15) can be rephrased equivalently in terms of the fluid-structure operator $A_{\mathrm{fs}}$ from (6.17) and the formulae for $(\operatorname{Id}_{3}-\mathbb{P})v$ and $p$ from (6.14). It follows from a combination of the liquid crystals part and the structure part as seen in Subsection 6.1.
Proposition 6.6.
Let $(f_{1},f_{2},f_{3},f_{4})\in\mathrm{X}$ and $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$. Then $(v,Q,\ell,\omega,\pi)$ with $v\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}$, $Q\in\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})$, $(\ell,\omega)\in\mathbb{C}^{3}\times\mathbb{C}^{3}$ and $\pi\in\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ is a solution to (6.15) if and only if
$\left\{\begin{aligned} (\lambda\operatorname{Id}-A_{\mathrm{fs}}(\widehat{Q}))\begin{pmatrix}\mathbb{P}v\\ Q\\ \ell\\ \omega\end{pmatrix}&=\begin{pmatrix}\mathbb{P}f_{1}\\ f_{2}\\ K^{-1}f_{3}\\ K^{-1}f_{4}\end{pmatrix},\kern 5.0pt(\operatorname{Id}_{3}-\mathbb{P})v=(\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega)\kern 5.0pt\text{and}\kern 5.0pt\\ \pi&=N(\Delta\mathbb{P}v\cdot n)-\lambda N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n).\end{aligned}\right$.
After defining the fluid-structure operator, we verify its maximal $\mathrm{L}^{p}$-regularity.
Proposition 6.7.
[…]
Proof.
[…]
$A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})\coloneqq\begin{pmatrix}-A_{\xi}(\widehat{Q})&0\\ 0&0\end{pmatrix}\kern 5.0pt\text{and}\kern 5.0ptB_{\mathrm{fs}}\coloneqq\begin{pmatrix}0&B_{1}\\ B_{2}&B_{3}\end{pmatrix}\coloneqq\begin{pmatrix}0&0&-A_{2}\mathbb{P}D_{\mathrm{fl}}\\ 0&0&0\\ K^{-1}C_{1}&K^{-1}C_{2}&K^{-1}C_{3}\end{pmatrix}$
$A_{\xi}(\widehat{Q}) fro$m (6.8) and the Stokes operator $A_{2}$ on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$. Besides, the Dirichlet lifting operator $D_{\mathrm{fl}}$ was made precise in (6.11), and the operators $C_{1}$, $C_{2}$ and $C_{3}$ and the matrix $K$ are defined in (6.16).
[…]
$\left\|B_{1}(\ell,\omega)\right\|_{\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0})}=\|-A_{2}\mathbb{P}D_{\mathrm{fl}}(\ell,\omega)\|_{\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})}\leq C\cdot\|D_{\mathrm{fl}}(\ell,\omega)\|_{\mathrm{H}^{2}(\mathcal{F}_{0})}\leq C\cdot\|(v,Q,\ell,\omega)\|_{\mathrm{X}}.$
This shows that $B_{1}\in\mathcal{L}(\mathbb{C}^{3}\times\mathbb{C}^{3},\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3}))$, so $B_{1}$ is a bounded perturbation. Moreover, we recall from the proof of [27, Theorem 3.11] that the operators $C_{1}$ and $C_{3}$ are relatively bounded with relative bound zero. For $C_{2}$, we invoke classical properties of the trace, interpolation and Young’s inequality to deduce that for any $\delta>0$, and with $\varepsilon>0$ small and $\theta\in(0,1)$, it holds that
$|C_{2}Q|\leq C\cdot\|Q\|_{\mathrm{H}^{\nicefrac{{5}}{{2}}+\varepsilon}(\mathcal{F}_{0})}\leq C\cdot\|Q\|_{\mathrm{H}^{1}(\mathcal{F}_{0})}^{\theta}\cdot\|Q\|_{\mathrm{H}^{3}(\mathcal{F}_{0})}^{1-\theta}\leq\delta\cdot\|Q\|_{\mathrm{H}^{3}(\mathcal{F}_{0})}+C(\delta)\cdot\|Q\|_{\mathrm{H}^{1}(\mathcal{F}_{0})}.$
[…]
6.3. Proof of Theorem 6.1
Theorem 6.1(a) directly follows from Subsection 6.2, while (b) relies on a refined spectral analysis.
Proof of Theorem 6.1(b).
For simplicity, we write $A_{\mathrm{fs}}\coloneqq A_{\mathrm{fs}}(0)$ in the sequel. It can be shown that the shift $\mu\geq 0$ in Subsection 6.2 can be be chosen equal to $s(A_{\mathrm{fs}})\coloneqq\sup\{\operatorname{Re}\lambda:\lambda\in\sigma(A_{\mathrm{fs}})\}$ of $A_{\mathrm{fs}}$, see for instance [35, Cor. 3.5.3]. Hence, to remove the shift in Subsection 6.2, and thus to obtain the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of $A_{\mathrm{fs}}$ on $\mathrm{X}$, we need to prove that the spectral bound satisfies $s(A_{\mathrm{fs}})<0$.
We start by noting that $A_{\mathrm{fs}}$ has a compact resolvent by compactness of the embedding $\mathrm{D}(A_{\mathrm{fs}})\hookrightarrow\mathrm{X}$ thanks to the Rellich-Kondrachov theorem. As a result, $\sigma(A_{\mathrm{fs}})=\sigma_{\mathrm{p}}(A_{\mathrm{fs}})$, i. e., the spectrum of $A_{\mathrm{fs}}$ only consists of eigenvalues. In view of Subsection 6.2, we may thus investigate (6.15) with right-hand side zero. Indeed, if for $\lambda\in\mathbb{C}$, the only solution to (6.15) with $(f_{1},f_{2},f_{3},f_{4})=(0,0,0,0)\in\mathrm{X}$ is zero, then Subsection 6.2 implies that $\lambda\notin\sigma_{p}(A_{\mathrm{fs}})$ and hence $\lambda\in\rho(A_{\mathrm{fs}})$ by the preceding argument.
Let now $\lambda\in\mathbb{C}$ and $(v,Q,\ell,\omega,\pi)\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{C}^{3}\times\mathbb{C}^{3}\times\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ satisfy (6.15) with $f_{i}=0$ for $i=1,\dots,4$. Next, we test the equation satisfied by $v$ in (6.15) by $\overline{v}$, the equation for $Q$ by $\Delta\overline{Q}$, the equation for $\ell$ by $\overline{\ell}$ and the equation for $\omega$ by $\overline{\omega}$. Integrating by parts, invoking Section 6(b) to handle the liquid crystals part, observing that $S_{\xi}(0)(-\Delta+\operatorname{Id}_{3})Q=\nicefrac{{2\xi}}{{3}}(\Delta Q-Q)$, making use of the interface condition from (2.8), and recalling $\mathrm{T}(v,\pi,Q)$ from (6.4), we derive that
$0$ $=\lambda\left(\|v\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\|\nabla Q\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+m_{\mathcal{S}}|\ell|^{2}+J_{0}\omega\cdot\overline{\omega}\right)$
$\quad+2\|\mathbb{D}(v)\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\|Q\|_{\mathrm{H}^{2}(\mathcal{F}_{0})}^{2}+\|\nabla Q\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\mathrm{i}\operatorname{Im}(2\langle\nabla v,S_{\xi}(0)(-\Delta+\operatorname{Id}_{3})Q\rangle_{\mathrm{L}^{2}(\mathcal{F}_{0})}.$
When considering the real part in the above equation, we obtain
$\operatorname{Re}\lambda\left(\|v\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\|\nabla Q\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+m_{\mathcal{S}}|\ell|^{2}\right)+\operatorname{Re}(\lambda J_{0}\omega\cdot\overline{\omega})+2\|\mathbb{D}(v)\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\|Q\|_{\mathrm{H}^{2}(\mathcal{F}_{0})}^{2}+\|\nabla Q\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}=0.$
It follows that $\operatorname{Re}\lambda\leq 0$. Moreover, if $\operatorname{Re}\lambda=0$, then $\mathbb{D}(v)=0$ and $Q=0$ on $\mathcal{F}_{0}$, so $v$ is especially constant. Due to $v=0$ on $\partial\mathcal{O}$, we infer that $v=0$ on $\mathcal{F}_{0}$. The interface condition $v=\ell+\omega\times y$ on $\partial\mathcal{S}_{0}$ then yields that $\ell=\omega=0$. From (6.15), it follows that $\nabla\pi=0$, and $\pi\in\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ rules out the constants, so $\pi=0$ and thus $(v,Q,\ell,\omega,\pi)=(0,0,0,0,0)$. The above arguments imply that $s(A_{\mathrm{fs}})<0$, resulting in the sectoriality and then also the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of $A_{\mathrm{fs}}$ on $\mathrm{X}$ even without shift.
In a similar manner as for Subsection 6.2, one can show that (6.5) admits an equivalent reformulation in terms of the fluid-structure operator $A_{\mathrm{fs}}$ from (6.17). Therefore, the first part of the assertion of Theorem 6.1(b) already follows, and the maximal regularity estimate is also implied in the case $\mu=0$.
By the sectoriality of $A_{\mathrm{fs}}$ with spectral angle $\phi_{-A_{\mathrm{fs}}}<\nicefrac{{\pi}}{{2}}$, the operator $A_{\mathrm{fs}}$ generates a bounded analytic semigroup $\left(\mathrm{e}^{tA_{\mathrm{fs}}}\right)_{t\geq 0}$, see e. g. [35, Thm. 3.3.2]. Thanks to $s(A_{\mathrm{fs}})<0$, this semigroup is even of negative exponential type. Thus, we may choose $\eta_{0}=-s(A_{\mathrm{fs}})>0$ and consider the operator $A_{\mathrm{fs}}+\eta$ for $\eta\in[0,\eta_{0})$, still generating a bounded analytic semigroup of negative exponential type thanks to the fact that $s(A_{\mathrm{fs}}+\eta)<0$, to deduce the estimate with exponential weights. ∎
7. Local strong well-posedness for large data
We use $z=(v,Q,\ell,\omega)$ again to shorten notation. First, we derive a reference solution to capture the initial data and to reduce the study to a problem with homogeneous initial values. The result below is a direct consequence of Theorem 6.1(a), because $z_{0}\in\mathrm{X}_{\gamma}$ implies $Q_{0}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$ for $p>4$.
Proposition 7.1.
Let $p>4$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, where $\mathrm{X}_{\gamma}$ has been introduced in (2.10), and consider $(f_{1},f_{2},f_{3},f_{4})=(0,0,0,0)$. Then for $\mu>\mu_{0}$, where $\mu_{0}\geq 0$ is as in Theorem 6.1(a), there is a unique solution $(z^{*},\pi^{*})\in\tilde{\mathbb{E}}_{\infty}$ to (6.5). In particular, we have $\|z^{*}-z_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\to 0$ as $T\to 0$.
[…]
(7.1) $\left\{\begin{aligned} \partial_{t}{\widehat{v}}+(\mu-\Delta){\widehat{v}}+\nabla{\widehat{\pi}}-\operatorname{div}S_{\xi}(Q_{0})(\Delta-\operatorname{Id}_{3})\widehat{Q}&=F_{1}(\widehat{z},\widehat{\pi}),\kern 5.0pt\operatorname{div}{\widehat{v}}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{\widehat{Q}}}+\tilde{S}_{\xi}(Q_{0})\nabla\widehat{v}+(\mu-(\Delta-\operatorname{Id}_{3})){\widehat{Q}}&=F_{2}(\widehat{z}),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\widehat{\ell}})^{\prime}+\mu\widehat{\ell}+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(\widehat{v},\widehat{\pi},\widehat{Q})n\,\mathrm{d}\Gamma&=F_{3}(\widehat{z},\widehat{\pi}),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ J_{0}({\widehat{\omega}})^{\prime}+\mu\widehat{\omega}+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(\widehat{v},\widehat{\pi},\widehat{Q})n\,\mathrm{d}\Gamma&=F_{4}(\widehat{z},\widehat{\pi}),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ {\widehat{v}}={\widehat{\ell}}+{\widehat{\omega}}\times y,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{S}_{0},\kern 5.0pt{\widehat{v}}=0,&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0pt\partial_{n}\widehat{Q}=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{F}_{0},\\ \widehat{v}(0)=0,\kern 5.0pt\widehat{Q}(0)=0,\kern 5.0pt\widehat{\ell}(0)&=0\kern 5.0pt\text{and}\kern 5.0pt\widehat{\omega}(0)=0.\end{aligned}\right$.
As the linear part of $\tau$ is already captured on the left-hand side of (7.1), we define the remaining part $\tau_{\mathrm{r}}$ by $\tau_{\mathrm{r}}(Q)\coloneqq\tau(Q)+\frac{2\xi}{3}(\Delta Q-Q)$. For the transformed terms as made precise in Section 5, and with
$\mathcal{T}(v,\pi,Q)\coloneqq\mathbb{O}(t)^{\top}\mathrm{T}(\mathbb{O}(t)v(t,y),\pi(t,y),Q(t,y))\mathbb{O}(t)$
for the rotation matrix $\mathbb{O}(t)$ related to (5.1), $\mathrm{T}(v,\pi,Q)$ as in (6.4) and $(z,\pi)=(\widehat{z}+z^{*},\widehat{\pi}+\pi^{*})$, we have
[…]
$\quad+\bigl(\mathcal{B}_{\tau_{\mathrm{h}}}(Q)+\mathcal{B}_{\sigma}(Q)-\operatorname{div}S_{\xi}(Q_{0})((\Delta-\operatorname{Id}_{3})Q)\bigr),$
[…]
$\quad+Q^{2}-\operatorname{tr}\bigl(Q^{2}\bigr)\nicefrac{{\operatorname{Id}_{3}}}{{3}}-\operatorname{tr}\bigl(Q^{2}\bigr)Q,$
[…]
We now fix $T_{0}$, $R_{0}>0$ and consider $T\in(0,T_{0}]$ and $R\in(0,R_{0}]$. Recalling $\mathbb{E}_{T}$, $\mathbb{E}_{T}^{\pi}$ and $\tilde{\mathbb{E}}_{T}$ from (6.2) and (6.3), we introduce the set $\mathcal{K}_{T}^{R}$ together with the solution map $\Phi_{T}^{R}$ defined by
[…]
Next, we address the nonlinear estimates. For $T\in(0,T_{0}]$, $R\in(0,R_{0}]$, let $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$. With $(z^{*},\pi^{*})$ from Section 7, we set $(z,\pi)\coloneqq(\tilde{v}+v^{*},\tilde{Q}+Q^{*},\tilde{\ell}+\ell^{*},\tilde{\omega}+\omega^{*},\tilde{\pi}+\pi^{*})$. As observed in [9, Remark 5.2], for $C_{T}^{*}\coloneqq\|z^{*}\|_{\mathbb{E}_{T}}+\|\pi^{*}\|_{\mathbb{E}_{T}^{\pi}}$, it is valid that $C_{T}^{*}\to 0$ as $T\to 0$. By construction, it also follows that
(7.3) $\|(z,\pi)\|_{\tilde{\mathbb{E}}_{T}}\leq R+C_{T}^{*}.$
We also set $C_{0}\coloneqq\|z_{0}\|_{\mathrm{X}_{\gamma}}=\|(v_{0},Q_{0},\ell_{0},\omega_{0})\|_{\mathrm{X}_{\gamma}}$. At this stage, we invoke the following well-known lemma, allowing for an estimate of $\sup_{t\in[0,T]}\|z(t)\|_{\mathrm{X}_{\gamma}}$ by the norm of the initial data and the solution, where the constant is independent of $T$.
Lemma 7.2.
For $z\in\mathbb{E}_{T}$, there is $T$-independent $C>0$ with $\sup_{t\in[0,T]}\|z(t)\|_{\mathrm{X}_{\gamma}}\leq C(\|z(0)\|_{\mathrm{X}_{\gamma}}+\|z\|_{\mathbb{E}_{T}})$. In particular, for $z^{*}$ from Section 7, we have $\|z^{*}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\leq C(C_{0}+C_{T}^{*})$.
In the sequel, we derive useful embeddings. The assertion of Section 7(a) is a consequence of the mixed derivative theorem, see e. g. [35, Corollary 4.5.10], and Sobolev embeddings. For part (b), we use [4, Theorem III.4.10.2] joint with interpolation theory and embeddings for the Besov spaces, see for example [44, Theorem 4.6.1].
Lemma 7.3.
For $p>4$ as well as $0<T\leq\infty$, recall $\mathbb{E}_{T}$ and $\mathrm{X}_{\gamma}$ from (6.2) and (2.10), respectively.
  1. (a)
    For every $\theta\in(0,1)$, we find that
    $\mathbb{E}_{T}\hookrightarrow\mathrm{H}^{\theta,p}\bigl(0,T;\mathrm{H}^{2(1-\theta)}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{2(1-\theta)+1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}\bigr)\kern 5.0pt\text{and}\kern 5.0pt\mathbb{E}_{T}^{Q}\hookrightarrow\mathrm{L}^{\infty}\bigl(0,T;\mathrm{H}^{2}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\bigr).$
  2. (b)
    It holds that $\mathbb{E}_{T}\hookrightarrow\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})\hookrightarrow\mathrm{BUC}\bigl([0,T];\mathrm{B}_{2p}^{2-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0})^{3}\times\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{6}\bigr)$, implying $\mathbb{E}_{T}\hookrightarrow\mathrm{BUC}\bigl([0,T];\mathrm{C}(\overline{\mathcal{F}_{0}})^{3}\times\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})\times\mathbb{R}^{6}\bigr)$. For $T=\infty$, $\mathrm{BUC}([0,T])$ is replaced by $\mathrm{L}^{\infty}(0,\infty)$.
The embedding constants can be chosen $T$-independent if the functions have homogeneous initial values.
Given the transformed body velocities, it is possible to successively deduce the matrix $\mathbb{O}$, the original body velocities $h^{\prime}$ and $\Omega$ as well as the diffeomorphisms $X$ and $Y$ as described in the context of the transformation to the fixed domain in Section 5. For more details, see [9, Remark 6.1].
In the sequel, we discuss useful estimates of $X$, $Y$ and related terms. For $(\ell_{i},\omega_{i})\in\mathrm{W}^{1,p}(0,T)^{6}$, the subscript $i$ is used to denote objects associated to $(\ell_{i},\omega_{i})$. Besides, we employ $\|\cdot\|_{\infty}$ and $\|\cdot\|_{\infty,\infty}$ for the $\mathrm{L}^{\infty}$-norm in time and time and space, respectively. The assertions of Section 7(a) and (b) follow upon modifying the arguments in [19, Section 6.1]. For the estimates of the second derivatives of $X$ and $Y$, note that the diffeomorphisms associated to $\ell=\omega=0$ are the identity, so their second derivatives vanish. Section 7(b) and Section 7 then imply $\|\ell_{i}\|_{\infty}+\|\omega_{i}\|_{\infty}\leq C(R+C_{0}+C_{T}^{*})$. For a proof of Section 7(c), we refer to [19, Lemma 6.5], while Section 7(d) can be deduced from (c).
Lemma 7.4.
Consider $(\ell,\omega)$, $(\ell_{1},\omega_{1})$, $(\ell_{2},\omega_{2})\in\mathrm{W}^{1,p}(0,T)^{6}$ as well as the associated diffeomorphisms $X$, $X_{1}$, $X_{2}$ and $Y$, $Y_{1}$ and $Y_{2}$.
  1. (a)
    For $i=1,2$, we have $X_{i}$, $Y_{i}\in\mathrm{C}^{1}\left(0,T;\mathrm{C}^{\infty}(\mathbb{R}^{3})^{3}\right)$, with $\|\partial^{\alpha}X_{i}\|_{\infty,\infty}+\|\partial^{\alpha}Y_{i}\|_{\infty,\infty}\leq C$ and
    $\|\partial^{\beta}(X_{1}-X_{2})\|_{\infty,\infty}+\|\partial^{\beta}(Y_{1}-Y_{2})\|_{\infty,\infty}\leq CT(\|\ell_{1}-\ell_{2}\|_{\infty}+\|\omega_{1}-\omega_{2}\|_{\infty})$
    for all multi-indices $\alpha$, $\beta$ with $1\leq|\alpha|\leq 3$ and $0\leq|\beta|\leq 3$. For $j,k,l\in\{1,2,3\}$, we have
    $\|\partial_{j}\partial_{k}(X_{i})_{l}\|_{\infty,\infty}$ $\leq CT(R+C_{0}+C_{T}^{*}),$ $\|\partial_{j}\partial_{k}(X_{i})_{l}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}\leq CT(R+C_{0}+C_{T}^{*}),$
    $\|\partial_{j}\partial_{k}(Y_{i})_{l}\|_{\infty,\infty}$ $\leq CT(R+C_{0}+C_{T}^{*})\kern 5.0pt\text{and}\kern 5.0pt$ $\|\partial_{j}\partial_{k}(Y_{i})_{l}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}\leq CT(R+C_{0}+C_{T}^{*}).$
  2. (b)
    The matrix $\mathbb{O}_{i}$ from Section 5 fulfills
    $\|\mathbb{O}_{1}-\mathbb{O}_{2}\|_{\infty}\leq CT\cdot\|M_{1}-M_{2}\|_{\infty}\leq CT\cdot\|\omega_{1}-\omega_{2}\|_{\infty}\kern 5.0pt\text{and}\kern 5.0pt\|\mathbb{O}_{i}\|_{\infty}+\|\mathbb{O}_{i}^{\top}\|_{\infty}\leq C.$
  3. (c)
    For every $\alpha$ with $0\leq|\alpha|\leq 1$, the covariant and contravariant tensor and the Christoffel symbol associated to $X$, $X_{1}$, $X_{2}$ and $Y$, $Y_{1}$, $Y_{2}$ fulfill $\|\partial^{\alpha}g^{ij}\|_{\infty,\infty}+\|\partial^{\alpha}g_{ij}\|_{\infty,\infty}+\|\partial^{\alpha}\Gamma_{jk}^{i}\|_{\infty,\infty}\leq C$ and
    $\|\partial^{\alpha}((g_{1})^{ij}-(g_{2})^{ij})\|_{\infty,\infty}+\|\partial^{\alpha}((g_{1})_{ij}-(g_{2})_{ij})\|_{\infty,\infty}+\|\partial^{\alpha}((\Gamma_{1})_{jk}^{i}-(\Gamma_{2})_{jk}^{i})\|_{\infty,\infty}\leq CT\cdot\|(\tilde{\ell}_{1}-\tilde{\ell}_{2},\tilde{\omega}_{1}-\tilde{\omega}_{2})\|_{\infty}.$
  4. (d)
    For $i,j\in\{1,2,3\}$, we have $\|g^{ij}-\delta_{ij}\|_{\infty,\infty}$, $\|\partial_{i}X_{j}-\delta_{ij}\|_{\infty,\infty}$, $\|\partial_{i}Y_{j}-\delta_{ij}\|_{\infty,\infty}\leq CT(R+C_{0}+C_{T}^{*})$. The estimates by $CT(R+C_{0}+C_{T}^{*})$ remain valid when $\|\cdot\|_{\infty,\infty}$ is replaced by $\|\cdot\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}$.
Below, we collect estimates of some terms from (7.2). Here terms associated to $z_{i}$ are indicated by (i). The estimates of the first four terms in Section 7(a) can be deduced from [19, Lemma 6.6], while the last two terms in (a) can be handled similarly as in [9, (6.8) and Proposition 6.6] upon invoking the estimates in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))$ from Section 7 and $\|\tilde{Q}\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}\leq T^{\nicefrac{{1}}{{p}}}\cdot\|\tilde{Q}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}\leq CT^{\nicefrac{{1}}{{p}}}\cdot\|\tilde{z}\|_{\mathbb{E}_{T}}$ for $\tilde{Q}\in\prescript{}{0}{\mathbb{E}_{T}^{Q}}$ by Section 7(a). For the estimates in Section 7(b) concerning the Cauchy stress tensor part, we refer to [19, Lemma 6.6], while the estimate of the remaining part can be derived from mapping properties of $\mathcal{J}$ defined below, the estimate $\|\mathbb{O}-\operatorname{Id}_{3}\|_{\infty}\leq CT$, following in turn from Section 7(b), and a suitable expansion of the difference.
Lemma 7.5.
Let $p>4$, recall $(z^{*},\pi^{*})$ from Section 7, and consider $(z,\pi)=(\tilde{z}+z^{*},\tilde{\pi}+\pi^{*})$ for $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$. Then there exists $C(R,T)>0$, with $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, such that
  1. (a)
    $\|(\mathcal{L}_{1}-\Delta)v\|_{\mathbb{F}_{T}^{v}}$, $\|\mathcal{M}v\|_{\mathbb{F}_{T}^{v}}$, $\|\mathcal{N}(v)\|_{\mathbb{F}_{T}^{v}}$, $\|(\mathcal{G}-\nabla)\pi\|_{\mathbb{F}_{T}^{v}}$, $\|(\mathcal{L}_{2}-\Delta)Q\|_{\mathbb{F}_{T}^{Q}}$, $\|(\partial_{t}Y\cdot\nabla)Q\|_{\mathbb{F}_{T}^{Q}}\leq C(R,T)$.
  2. (b)
    For $\varepsilon>0$, define $\mathcal{J}\colon\mathrm{H}^{\varepsilon+\nicefrac{{1}}{{2}}}(\mathcal{F}_{0})^{3\times 3}\to\mathbb{R}^{6}$ by $\mathcal{J}(h)\coloneqq(\int_{\partial\mathcal{S}_{0}}hn\,\mathrm{d}\Gamma,\int_{\partial\mathcal{S}_{0}}y\times hn\,\mathrm{d}\Gamma)^{\top}$. Then there is $C>0$ with $\|\mathcal{J}((\mathrm{T}-\mathcal{T})(v,\pi,Q))\|_{p}\leq CT\cdot\|z\|_{\mathbb{E}_{T}}$.
The lemma below on Lipschitz estimates of the non-transformed terms follows from [22, Lemma 6.3].
Lemma 7.6.
Let $p>4$ and $z_{1}$, $z_{2}\in\mathrm{X}_{\gamma}$, with $\mathrm{X}_{\gamma}$ as defined in (2.10), and recall $A_{\xi}$ from (6.8). Then there exists $C>0$ such that $\|A_{\xi}(Q_{1})-A_{\xi}(Q_{2})\|_{\mathcal{L}(\mathrm{H}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{3}(\mathcal{F}_{0}),\mathrm{L}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0}))}\leq C\cdot\|z_{1}-z_{2}\|_{\mathrm{X}_{\gamma}}$.
As a preparation of the estimates of $F_{1}$, we now address the second line of $F_{1}$ as made precise in (7.2).
Lemma 7.7.
Consider $p>4$, and for the reference solution $(z^{*},\pi^{*})$ from Section 7 and $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$, let $(z,\pi)=(\tilde{z}+z^{*},\tilde{\pi}+\pi^{*})$. Then there exists $C(R,T)>0$ such that $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, and with $\|\mathcal{B}_{\tau_{\mathrm{h}}}(\tilde{Q}+Q^{*})+\mathcal{B}_{\sigma}(\tilde{Q}+Q^{*})-\operatorname{div}S_{\xi}(Q_{0})(\Delta-\operatorname{Id}_{3})(\tilde{Q}+Q^{*})\|_{\mathbb{F}_{T}^{v}}\leq C(R,T)$.
Proof.
The idea is to add and subtract terms, leading to two estimates which only concern the non-transformed terms and an estimate of the difference of transformed and non-transformed terms. From Section 5, we recall $B_{\tau_{\mathrm{h}}}(\tilde{Q}+Q^{*})(\tilde{Q}+Q^{*})+B_{\sigma}(\tilde{Q}+Q^{*})(\tilde{Q}+Q^{*})=\operatorname{div}S_{\xi}(\tilde{Q}+Q^{*})(\Delta-\operatorname{Id}_{3})(\tilde{Q}+Q^{*})$. With this identity, we expand the difference to be estimated as
$\mathrm{I}+\mathrm{II}+\mathrm{III}$ $\coloneqq\mathcal{B}_{\tau_{\mathrm{h}}}(\tilde{Q}+Q^{*})-B_{\tau_{\mathrm{h}}}(\tilde{Q}+Q^{*})(\tilde{Q}+Q^{*})+\mathcal{B}_{\sigma}(\tilde{Q}+Q^{*})-B_{\sigma}(\tilde{Q}+Q^{*})(\tilde{Q}+Q^{*})$
$\quad+\operatorname{div}S_{\xi}(\tilde{Q}+Q^{*})(\Delta-\operatorname{Id}_{3})(\tilde{Q}+Q^{*})-\operatorname{div}S_{\xi}(Q^{*})(\Delta-\operatorname{Id}_{3})(\tilde{Q}+Q^{*})$
$\quad+\operatorname{div}S_{\xi}(Q^{*})(\Delta-\operatorname{Id}_{3})(\tilde{Q}+Q^{*})-\operatorname{div}S_{\xi}(Q_{0})(\Delta-\operatorname{Id}_{3})(\tilde{Q}+Q^{*}).$
With regard to the terms $\mathrm{II}$ and $\mathrm{III}$, recall that $\tilde{z}$, $z^{*}\in\mathbb{E}_{T}$ by construction. Hence, Section 7(b) implies that $\tilde{z}(t)+z^{*}(t)$, $z^{*}(t)\in\mathrm{X}_{\gamma}$ for all $t\in[0,T]$, and $z_{0}\in\mathrm{X}_{\gamma}$ is valid by assumption. Thus, Section 7 and the estimate of $\|z\|_{\mathbb{E}_{T}}$ from (7.3) as well as Section 7(b) imply that
$\|\mathrm{II}\|_{\mathbb{F}_{T}^{v}}$ $\leq\Bigl(\int_{0}^{T}\bigl\|\bigl(A_{\xi}(\tilde{Q}(t)+Q^{*}(t))-A_{\xi}(Q^{*}(t))\bigr)(\tilde{v}(t)+v^{*}(t),\tilde{Q}(t)+Q^{*}(t))^{\top}\bigr\|_{\mathrm{L}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0})}^{p}\,\mathrm{d}t\Bigr)^{\nicefrac{{1}}{{p}}}$
$\leq C\Bigl(\int_{0}^{T}\|\tilde{z}(t)\|_{\mathrm{X}_{\gamma}}^{p}\cdot\|\tilde{z}(t)+z^{*}(t)\|_{\mathrm{X}_{1}}^{p}\,\mathrm{d}t\Bigr)^{\nicefrac{{1}}{{p}}}\leq C\cdot\|\tilde{z}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\cdot\|z\|_{\mathbb{E}_{T}}\leq CR(R+C_{T}^{*}).$
In a similar way, we find that $\|\mathrm{III}\|_{\mathbb{F}_{T}^{v}}\leq C(R+C_{T}^{*})\cdot\|z^{*}-z_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}$. From Section 7, we recall that $\|z^{*}-z_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}$ tends to zero as $T\to 0$.
It thus remains to estimate $\mathrm{I}$. We split this term into the part associated to $\tau_{\mathrm{h}}$ and $\sigma$. For brevity, we only elaborate on the estimate of $\mathrm{I}_{\sigma}\coloneqq\mathcal{B}_{\sigma}(\tilde{Q}+Q^{*})-B_{\sigma}(\tilde{Q}+Q^{*})(\tilde{Q}+Q^{*})$ and remark that handling the difference $\mathcal{B}_{\tau_{\mathrm{h}}}(\tilde{Q}+Q^{*})-B_{\tau_{\mathrm{h}}}(\tilde{Q}+Q^{*})(\tilde{Q}+Q^{*})$ is conceptually similar. In fact, the idea is to expand the terms to get differences which allow for estimates by a positive power of $T$ by Section 7.
With regard to $\mathrm{I}_{\sigma}$, recalling $B_{\sigma}(\tilde{Q}+Q^{*})(\tilde{Q}+Q^{*})$ and $\mathcal{B}_{\sigma}(\tilde{Q}+Q^{*})$ from (5.5) and (5.7), respectively, we remark that it suffices to consider the first two terms in the difference by symmetry. In the sequel, we concentrate on the second term, as it involves the highest derivatives. The first term in the difference $\mathrm{I}_{\sigma,1}$ can be treated in a similar fashion. Thus, we expand $\mathrm{I}_{\sigma,2}\coloneqq\mathcal{B}_{\sigma,12}(Q)Q-\sum_{j,k=1}^{3}Q_{ik}(\partial_{j}\Delta Q_{kj})$ as
$\mathrm{I}_{\sigma,2}$ $=\sum_{j,k=1}^{3}Q_{ik}\Biggl(\sum_{l=1}^{3}\biggl(\sum_{m=1}^{3}\Bigl(\sum_{n=1}^{3}(\partial_{n}\partial_{m}\partial_{l}Q_{kj})(\partial_{j}Y_{n}-\delta_{jn})g^{lm}+(\partial_{n}\partial_{m}\partial_{l}Q_{kj})\delta_{jn}(g^{lm}-\delta_{lm})\Bigr)$
$\qquad+(\partial_{m}\partial_{l}Q_{kj})(\partial_{j}g^{lm})+(\partial_{m}\partial_{l}Q_{kj})(\partial_{j}Y_{m})\Delta Y_{l}\biggr)+(\partial_{l}Q_{kj})(\partial_{j}\Delta Y_{l})\Biggr).$
The factor $Q_{ik}$ can be bounded in $\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\mathcal{F}_{0}))$, since the embeddings from Section 7(b) and the estimate of $\|z^{*}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}$ from Section 7 especially imply that
(7.4) $\|Q\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}\leq C(\|\tilde{z}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}+\|z^{*}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})})\leq C(R+C_{0}+C_{T}^{*}).$
Now, we derive estimates of second derivatives of $Q$. As $p>4$, there is $q>p$ with $0<\eta\coloneqq\nicefrac{{1}}{{p}}-\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$. From Section 7(a) and Sobolev embeddings, and for $\theta\in(\eta,\nicefrac{{1}}{{2}})$, it then follows that
$\|Q\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}\leq T^{\eta}\cdot\|\tilde{Q}\|_{\mathrm{L}^{q}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}+C_{T}^{*}\leq CT^{\eta}\cdot\|\tilde{Q}\|_{\mathrm{H}^{\theta,p}(0,T;\mathrm{H}^{2(1-\theta)+1}(\mathcal{F}_{0}))}+C_{T}^{*}\leq CT^{\eta}R+C_{T}^{*}.$
From (7.3), we deduce $\|\partial^{\alpha}Q_{kj}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{2}(\mathcal{F}_{0}))}\leq\|z\|_{\mathbb{E}_{T}}\leq R+C_{T}^{*}$ for all multi-indices $\alpha$ with $0\leq|\alpha|\leq 3$. By virtue of Section 7(d), the differences $\partial_{j}Y_{n}-\delta_{jn}$ and $g^{lm}-\delta_{lm}$ admit estimates by $CT(R+C_{0}+C_{T}^{*})$ in $\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\mathcal{F}_{0}))$. On the other hand, the boundedness of terms as $\partial_{j}g^{lm}$, $\partial_{j}Y_{m}$, $\Delta Y_{l}$ and $\partial_{j}\Delta Y_{l}$ follows from Section 7(c) and (a), while a decay estimate of the term $\partial_{l}Q_{kj}$ is a consequence of the above estimate of $\|Q\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{2}(\mathcal{F}_{0}))}$. In total, for some $C_{\mathrm{I}_{\sigma,2}}(R,T)>0$ with $C_{\mathrm{I}_{\sigma,2}}(R,T)\to 0$ as $R\to 0$ and $T\to 0$, we conclude $\|\mathrm{I}_{\sigma,2}\|_{\mathbb{F}_{T}^{v}}\leq C_{\mathrm{I}_{\sigma,2}}(R,T)$. In view of the above remarks, this completes the proof. ∎
The next proposition asserts the self-map and Lipschitz estimates of $F_{1}$. It is a consequence of Section 7, Section 7 and the observation that the estimates of the transformed lower order terms $\mathcal{B}_{\tau_{\mathrm{l}}}$ are conceptually similar to those in Section 7. The Lipschitz estimates follow in the same way as the self-map estimate, see also [19, Section 6.1] or [9, Section 6.2] for examples of such Lipschitz estimates.
Proposition 7.8.
[…]
$\|F_{1}(\tilde{z},\tilde{\pi})\|_{\mathbb{F}_{T}^{v}}\leq C_{1}(R,T)\kern 5.0pt\text{and}\kern 5.0pt\|F_{1}(\tilde{z}_{1},\tilde{\pi}_{1})-F_{1}(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\mathbb{F}_{T}^{v}}\leq L_{1}(R,T)\cdot\|(\tilde{z}_{1},\tilde{\pi}_{1})-(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}}.$
In order to estimate $F_{2}$, we also deal with the most difficult term in advance.
Lemma 7.9.
Let $p>4$, and for the reference solution $(z^{*},\pi^{*})$ from Section 7 and $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$, let $(z,\pi)=(\tilde{z}+z^{*},\tilde{\pi}+\pi^{*})$ and $(z_{i},\pi_{i})=(\tilde{z}_{i}+z^{*},\tilde{\pi}_{i}+\pi^{*})$. Then there is $C(R,T)>0$, with $C(R,T)\to 0$ as $R\to 0$ and $T\to 0$, such that $\|\mathcal{S}(\tilde{v}+v^{*},\tilde{Q}+Q^{*})+\tilde{S}_{\xi}(Q_{0})\nabla(\tilde{v}+v^{*})\|_{\mathbb{F}_{T}^{Q}}\leq C(R,T)$.
Proof.
As in the proof of Section 7, the idea here is to split the difference suitably. In fact, invoking that $S(\nabla v,Q)=\tilde{S}_{\xi}(Q)\nabla v$ from Section 3, we write
$\mathrm{i}+\mathrm{ii}+\mathrm{iii}$ $\coloneqq\bigl(\mathcal{S}(\tilde{v}+v^{*},\tilde{Q}+Q^{*})-S(\nabla(\tilde{v}+v^{*}),\tilde{Q}+Q^{*})\bigr)-\bigl(\tilde{S}_{\xi}(\tilde{Q}+Q^{*})\nabla(\tilde{v}+v^{*})-\tilde{S}_{\xi}(Q^{*})\nabla(\tilde{v}+v^{*})\bigr)$
$\quad-\bigl(\tilde{S}_{\xi}(Q^{*})\nabla(\tilde{v}+v^{*})-\tilde{S}_{\xi}(Q_{0})\nabla(\tilde{v}+v^{*})\bigr).$
From the proof of Section 7, we recall that $\tilde{z}(t)+z^{*}(t)$, $z^{*}(t)\in\mathrm{X}_{\gamma}$ for all $t\in[0,T]$, so Section 7 is applicable. As for the estimate of $\mathrm{II}$ and $\mathrm{III}$ in the proof of Section 7, it then follows that
$\|\mathrm{ii}\|_{\mathbb{F}_{T}^{Q}}=\bigl\|\bigl(\tilde{S}_{\xi}(\tilde{Q}+Q^{*})-\tilde{S}_{\xi}(Q^{*})\bigr)\nabla(\tilde{v}+v^{*})\bigr\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{1}(\mathcal{F}_{0}))}\leq C\cdot\|\tilde{z}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\cdot\|z\|_{\mathbb{E}_{T}}\leq CR(R+C_{T}^{*}),$
$\|\mathrm{iii}\|_{\mathbb{F}_{T}^{Q}}=\bigl\|\bigl(\tilde{S}_{\xi}(Q^{*})-\tilde{S}_{\xi}(Q_{0})\bigr)\nabla(\tilde{v}+v^{*})\bigr\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{1}(\mathcal{F}_{0}))}\leq C(R+C_{T}^{*})\cdot\|z^{*}-z_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}.$
The last factor tends to zero as $T\to 0$ by Section 7, showing the estimates for $\mathrm{ii}$ and $\mathrm{iii}$. It remains to deal with the term $\mathrm{i}$. To this end, recalling $\mathcal{D}$ from (5.9), we expand the difference $\mathcal{D}_{v}\coloneqq\mathcal{D}-\nabla v$ as
(7.5) $(\mathcal{D}_{v})_{ij}=\sum_{k=1}^{3}\sum_{l=1}^{3}(\partial_{l}\partial_{k}X_{i})(\partial_{j}Y_{l})v_{k}+\sum_{k=1}^{3}(\partial_{k}X_{i}-\delta_{ki})\sum_{l=1}^{3}(\partial_{l}v_{k})(\partial_{j}Y_{l})+\sum_{l=1}^{3}\partial_{l}v_{i}(\partial_{j}Y_{l}-\delta_{jl}).$
Making use of Hölder’s inequality together with Section 7 in order to estimate the second derivatives of $X$ and the differences $\partial_{k}X_{i}-\delta_{ki}$ and $\partial_{j}Y_{l}-\delta_{jl}$ by $T(R+C_{0}+C_{T}^{*})$, we infer that
(7.6) $\|\mathcal{D}_{v}\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{1}(\mathcal{F}_{0}))}\leq CT(R+C_{0}+C_{T}^{*})(R+C_{T}^{*}),$
and likewise for $\mathcal{D}_{v}^{\top}=\mathcal{D}^{\top}-\nabla v^{\top}$. We recall the respective terms from Section 5 and get
$\mathrm{i}=-\frac{1}{2}\bigl([Q,\mathcal{D}_{v}]-[Q,\mathcal{D}_{v}^{\top}]\bigr)+\xi\Bigl(\frac{1}{3}\bigl(\mathcal{D}_{v}+\mathcal{D}_{v}^{\top}\bigr)+\frac{1}{2}\bigl(\{Q,\mathcal{D}_{v}\}+\{Q,\mathcal{D}_{v}^{\top}\}\bigr)-2(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(Q\mathcal{D}_{v})\Bigr).$
Observe that every addend contains a factor of the form $\mathcal{D}-\nabla v$ or $\mathcal{D}^{\top}-\nabla v^{\top}$, so (7.6) can be used. For the terms involving products with $Q$, we employ Hölder’s inequality and invoke the estimate of $\|Q\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}$ from (7.4). In conclusion, we find that there exists $C_{\mathrm{i}}(R,T)$, with $C_{\mathrm{i}}(R,T)\to 0$ as $R\to 0$ and $T\to 0$, such that $\|\mathrm{i}\|_{\mathbb{F}_{T}^{Q}}\leq C_{\mathrm{i}}(R,T)$, finishing the proof. ∎
The proposition below asserts that $F_{2}$ admits the desired estimates.
Proposition 7.10.
[…]
$\|F_{2}(\tilde{z})\|_{\mathbb{F}_{T}^{Q}}\leq C_{2}(R,T)\kern 5.0pt\text{and}\kern 5.0pt\|F_{2}(\tilde{z}_{1})-F_{2}(\tilde{z}_{2})\|_{\mathbb{F}_{T}^{Q}}\leq L_{2}(R,T)\cdot\|\tilde{z}_{1}-\tilde{z}_{2}\|_{\mathbb{E}_{T}}.$
Proof.
In view of Section 7, Section 7 and the fact that the estimates of $\mu Q$ are simple, we still need to estimate $((\tilde{v}+v^{*})\cdot\nabla)(\tilde{Q}+Q^{*})$, $(\tilde{Q}+Q^{*})^{2}-\operatorname{tr}((\tilde{Q}+Q^{*})^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}}$ and $\operatorname{tr}((\tilde{Q}+Q^{*})^{2})(\tilde{Q}+Q^{*})$.
First, we discuss estimates of powers of $Q$. For $m\geq 2$, recalling the estimate of $\|Q\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}$ from (7.4), and using that Section 7(b) implies $\|Q\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{1}(\mathcal{F}_{0}))}\leq C_{T}^{*}+T^{\nicefrac{{1}}{{p}}}R$, we obtain
(7.7) $\|Q^{m}\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{1}(\mathcal{F}_{0}))}\leq\|Q\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}^{m-1}\cdot\|Q\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{1}(\mathcal{F}_{0}))}\leq C(C_{0}+R+C_{T}^{*})^{m-1}(C_{T}^{*}+T^{\nicefrac{{1}}{{p}}}R).$
As $\operatorname{tr}((\tilde{Q}+Q^{*})^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}}$ can be handled similarly, we get $\|Q^{2}-\operatorname{tr}(Q^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}}\|_{\mathbb{F}_{T}^{Q}}\leq C(C_{0}+R+C_{T}^{*})(C_{T}^{*}+T^{\nicefrac{{1}}{{p}}}R)$.
Concerning the estimate of $(v\cdot\nabla)Q$, we note that $\mathrm{H}^{\theta,p}\bigl(0,T;\mathrm{H}^{2(1-\theta)}(\mathcal{F}_{0})\bigr)\hookrightarrow\mathrm{L}^{q}\bigl(0,T;\mathrm{W}^{1,2s}(\mathcal{F}_{0})\bigr)$ holds for $q>p$ and $\theta\in(\nicefrac{{1}}{{p}}-\nicefrac{{1}}{{q}},\nicefrac{{3}}{{4s}}-\nicefrac{{1}}{{4}})$, requiring $s<3$, or, equivalently, $s^{\prime}>\nicefrac{{3}}{{2}}$ for $\nicefrac{{1}}{{s}}+\nicefrac{{1}}{{s^{\prime}}}=1$. For such $s$, Section 7(a) yields $\|v\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,2s}(\mathcal{F}_{0}))}\leq T^{\eta}\cdot\|\tilde{v}\|_{\mathrm{L}^{q}(0,T;\mathrm{W}^{1,2s}(\mathcal{F}_{0}))}+\|z^{*}\|_{\mathbb{E}_{T}}\leq CT^{\eta}R+C_{T}^{*}$ for $\eta\coloneqq\nicefrac{{1}}{{p}}-\nicefrac{{1}}{{q}}$. On the other hand, thanks to $p>4$, there is $s^{\prime}>\nicefrac{{3}}{{2}}$ such that $\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0})\hookrightarrow\mathrm{W}^{2,2s^{\prime}}(\mathcal{F}_{0})$. By Section 7(b), and with $\|z\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\leq C_{0}+R+C_{T}^{*}$, we get $\|\nabla Q\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2s^{\prime}}(\mathcal{F}_{0}))}\leq C(C_{0}+R+C_{T}^{*})$. Hölder’s inequality and a combination of the above estimates then results in
$\|(v\cdot\nabla)Q\|_{\mathbb{F}_{T}^{Q}}\leq\|v\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,2s}(\mathcal{F}_{0}))}\cdot\|\nabla Q\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2s^{\prime}}(\mathcal{F}_{0}))}\leq C(C_{0}+R+C_{T}^{*})(T^{\eta}R+C_{T}^{*}).$
Let us observe that the Lipschitz estimate can be obtained in a similar way. ∎
Finally, we estimate the terms $F_{3}$ and $F_{4}$. The complexity of the stress tensor and the present anisotropic Hilbert space setting renders renders this task challenging.
Proposition 7.11.
[…]
$\|F_{3}(\tilde{z},\tilde{\pi})\|_{\mathbb{F}_{T}^{\ell}}$ $\leq C_{3}(R,T),\kern 5.0pt\|F_{3}(\tilde{z}_{1},\tilde{\pi}_{1})-F_{3}(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\mathbb{F}_{T}^{\ell}}\leq L_{3}(R,T)\cdot\|(\tilde{z}_{1},\tilde{\pi}_{1})-(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}}\kern 5.0pt\text{as well as}\kern 5.0pt$
$\|F_{4}(\tilde{z},\tilde{\pi})\|_{\mathbb{F}_{T}^{\omega}}$ $\leq C_{4}(R,T),\kern 5.0pt\|F_{4}(\tilde{z}_{1},\tilde{\pi}_{1})-F_{4}(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\mathbb{F}_{T}^{\omega}}\leq L_{4}(R,T)\cdot\|(\tilde{z}_{1},\tilde{\pi}_{1})-(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}}.$
Proof.
Note that the estimates of $\omega\times\ell$, $\omega\times(J_{0}\omega)$, $\mu\ell$ and $\mu\omega$ are straightforward, while the estimates of the Cauchy part of the stress tensor follow from Section 7(b). Hence, it remains to handle the surface integral terms involving $\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)$, where the most difficult terms are $\nabla Q\odot\nabla Q$, $2\xi Q\operatorname{tr}(Q\Delta Q)$ and $2\xi Q\operatorname{tr}(Q^{2})\operatorname{tr}(Q^{2})$. The study of the other terms can be reduced to the latter ones.
The continuity of the trace from $\mathrm{H}^{\nicefrac{{1}}{{2}}+\varepsilon}(\mathcal{F}_{0})$ to $\mathrm{L}^{2}(\partial\mathcal{S}_{0})$, $\varepsilon>0$, the bound of $\|\mathbb{O}\|_{\infty}$ by Section 7(b) and the embedding from Section 7(b) together with $\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0})\hookrightarrow\mathrm{H}^{\nicefrac{{3}}{{2}}+\varepsilon}(\mathcal{F}_{0})$ first yield
$\left\|\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}\nabla Q\odot\nabla Q\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}$ $\leq C\cdot\|\tilde{Q}+Q^{*}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{H}^{\nicefrac{{3}}{{2}}+\varepsilon}(\mathcal{F}_{0}))}\cdot\|\tilde{Q}+Q^{*}\|_{\mathrm{L}^{p}(0,T;\mathrm{H}^{\nicefrac{{3}}{{2}}+\varepsilon}(\mathcal{F}_{0}))}$
$\leq C\left(\|\tilde{z}\|_{\mathbb{E}_{T}}+\|z^{*}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\right)\left(T^{\nicefrac{{1}}{{p}}}\cdot\|\tilde{Q}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{H}^{\nicefrac{{3}}{{2}}+\varepsilon}(\mathcal{F}_{0}))}+\|z^{*}\|_{\mathbb{E}_{T}}\right)$
$\leq C(R+C_{T}^{*}+C_{0})(T^{\nicefrac{{1}}{{p}}}R+C_{T}^{*}).$
We proceed with the estimate of $2\xi Q\operatorname{tr}(Q\Delta Q)$, and as $\xi\in\mathbb{R}$ is fixed, we omit it in the sequel. Similarly as above, using $\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0})\hookrightarrow\mathrm{W}^{\nicefrac{{1}}{{4}}+\varepsilon,4}(\mathcal{F}_{0})$ for the first two factors, and employing Section 7(a) joint with $\mathrm{H}^{\theta,p}\bigl(0,T;\mathrm{H}^{1+2(1-\theta)}(\mathcal{F}_{0})\bigr)\hookrightarrow\mathrm{L}^{q}\bigl(0,T;\mathrm{H}^{\nicefrac{{5}}{{2}}+\varepsilon}(\mathcal{F}_{0})\bigr)$ for $q>p$ and $\theta\in(\nicefrac{{1}}{{p}}-\nicefrac{{1}}{{q}},\nicefrac{{1}}{{4}})$, we get
$\left\|\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}Q\operatorname{tr}(Q\Delta Q)\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}$ $\leq C(R+C_{T}^{*}+C_{0})^{2}\left(T^{\eta}\cdot\|\tilde{Q}\|_{\mathrm{L}^{q}(0,T;\mathrm{H}^{\nicefrac{{5}}{{2}}+\varepsilon}(\mathcal{F}_{0}))}+\|z^{*}\|_{\mathbb{E}_{T}}\right)$
$\leq(R+C_{T}^{*}+C_{0})^{2}(T^{\eta}R+C_{T}^{*}),$
where $\eta=\nicefrac{{1}}{{p}}-\nicefrac{{1}}{{q}}>0$. The last term is $2\xi Q\operatorname{tr}(Q^{2})\operatorname{tr}(Q^{2})$, and we observe that $Q\operatorname{tr}(Q^{2})=Q|Q|_{2}^{4}$ for the Hilbert-Schmidt norm $|\cdot|_{2}$. The resulting first four factors can be treated as above upon noting that $\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0})\hookrightarrow\mathrm{W}^{\nicefrac{{1}}{{8}}+\varepsilon,8}(\mathcal{F}_{0})$. The last factor can be handled as in the preceding estimate, so
$\left\|\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}Q\operatorname{tr}(Q^{2})\operatorname{tr}(Q^{2})\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C(R+C_{T}^{*}+C_{0})^{4}(T^{\eta}R+C_{T}^{*}).$
Similar arguments yield the Lipschitz estimate of $F_{3}$ by the multilinear structure of the terms. Finally, the estimates of the surface integrals for $F_{4}$ are completely analogous to those of $F_{3}$. ∎
We are now in the position to show the local strong well-posedness for large data.
Proof of Theorem 4.1.
We start with the proof of Theorem 5.1, so let $(\tilde{z},\tilde{\pi})\in\mathcal{K}_{T}^{R}$. Theorem 6.1(a) applied to (7.1) together with Section 7, Section 7 and Section 7 results in
$\|\Phi_{T}^{R}(\tilde{z},\tilde{\pi})\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}\cdot\|(F_{1}(\tilde{z},\tilde{\pi}),F_{2}(\tilde{z}),F_{3}(\tilde{z},\tilde{\pi}),F_{4}(\tilde{z},\tilde{\pi}))\|_{\mathbb{F}_{T}}\leq C_{\mathrm{MR}}C(R,T).$
For given $R>0$, we get $C(R,T)<\nicefrac{{R}}{{2C_{\mathrm{MR}}}}$ provided $T>0$ is sufficiently small. In the same way,
$\|\Phi_{T}^{R}(\tilde{z}_{1},\tilde{\pi}_{1})-\Phi_{T}^{R}(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}L(R,T)\cdot\|(\tilde{z}_{1},\tilde{\pi}_{1})-(\tilde{z}_{2},\tilde{\pi}_{2})\|_{\tilde{\mathbb{E}}_{T}},\kern 5.0pt\text{for}\kern 5.0pt(\tilde{z}_{1},\tilde{\pi}_{1}),\,(\tilde{z}_{2},\tilde{\pi}_{2})\in\mathcal{K}_{T}^{R}.$
As $C_{\mathrm{MR}}>0$ is $T$-independent by virtue of the homogeneous initial values, we may choose $R>0$ and then also $T>0$ sufficiently small such that $C_{\mathrm{MR}}C(R,T)\leq\nicefrac{{R}}{{2}}$ and $C_{\mathrm{MR}}L(R,T)\leq\nicefrac{{1}}{{2}}$. In other words, the solution map $\Phi_{T}^{R}$ to the linearized problem (7.1) is a self-map and contraction, giving rise to a unique fixed point $(\widehat{z},\widehat{\pi})\in\prescript{}{0}{\mathbb{E}_{T}}\times\mathbb{E}_{T}^{\pi}$. An addition of the reference solution $(z^{*},\pi^{*})$ from Section 7 leads to the unique solution $(z,\pi)=(\widehat{z}+z^{*},\widehat{\pi}+\pi^{*})\in\tilde{\mathbb{E}}_{T}$ to (5.11)–(5.13). The proof is completed upon noting that $(z,\pi)\in\tilde{\mathbb{E}}_{T}$ means that $(v,Q,\ell,\omega,\pi)$ is in the regularity class as stated in Theorem 5.1.
Finally, we deduce Theorem 4.1 from Theorem 5.1. Thus, consider $(\ell,\omega)\in\mathrm{W}^{1,p}(0,T)^{6}$ resulting from Theorem 5.1. We recover $h^{\prime}$, $\Omega$ and $X$ from there. The backward change of variables from Section 5 yields the assertion of Theorem 4.1, where the uniqueness follows from the uniqueness of the diffeomorphism. ∎
8. Global strong well-posedness for small data
The purpose of this section is to prove Theorem 4.2 on the global strong well-posedness of the interaction problem (2.5)–(2.9) for small initial data. To this end, we set up a fixed point argument based on the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of the fluid-structure operator, and we derive new decay estimates of the coordinate transform in order to estimate the nonlinear terms.
In view of Theorem 6.1(b), we consider here the linearization at zero, without shift, and without invoking a reference solution thanks to the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity, i. e., we take into account
(8.1) $\left\{\begin{aligned} \partial_{t}{v}-\Delta{v}+\nabla\pi-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q&=G_{1}(z,\pi),\kern 5.0pt\operatorname{div}{v}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(0)\nabla v-(\Delta-\operatorname{Id}_{3}){Q}&=G_{2}(z),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\ell})^{\prime}+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=G_{3}(z,\pi),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ J_{0}({\omega})^{\prime}+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=G_{4}(z,\pi),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ {v}={\ell}+{\omega}\times y,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{S}_{0},\kern 5.0pt{v}&=0,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0pt\partial_{n}Q=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\kern 5.0ptQ(0)=Q_{0},\kern 5.0pt\ell(0)&=\ell_{0}\kern 5.0pt\text{and}\kern 5.0pt\omega(0)=\omega_{0}.\end{aligned}\right$.
For the transformed terms as made precise in Section 5, the terms on the right-hand side in (8.1) are
(8.2) $G_{1}(z,\pi)$ $\coloneqq(\mathcal{L}_{1}-\Delta)v-\mathcal{M}v-\mathcal{N}(v)-(\mathcal{G}-\nabla)\pi+\mathcal{B}_{\tau}(Q)+\mathcal{B}_{\sigma}(Q)-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q,$
$G_{2}(z)$ $\coloneqq(\mathcal{L}_{2}-\Delta)Q-(\partial_{t}Y\cdot\nabla)Q-(v\cdot\nabla)Q+\mathcal{S}(v,Q)+\tilde{S}_{\xi}(0)\nabla v+Q^{2}-\operatorname{tr}(Q^{2})(\nicefrac{{\operatorname{Id}_{3}}}{{3}}+Q),$
$G_{3}(z,\pi)$ $\coloneqq-m_{\mathcal{S}}\omega\times\ell+\int_{\partial\mathcal{S}_{0}}\bigl((\mathrm{T}-\mathcal{T})(v,\pi,Q)-\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)\bigr)n\,\mathrm{d}\Gamma\kern 5.0pt\text{and}\kern 5.0pt$
$G_{4}(z,\pi)$ $\coloneqq\omega\times(J_{0}\omega)+\int_{\partial\mathcal{S}_{0}}y\times\bigl((\mathrm{T}-\mathcal{T})(v,\pi,Q)-\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)\bigr)n\,\mathrm{d}\Gamma.$
For $\eta\in(0,\eta_{0})$, where $\eta_{0}>0$ has been introduced in Theorem 6.1(b), and $\tilde{\mathbb{E}}_{\infty}$ as made precise in (6.3), we define $\tilde{\mathcal{K}}$ to be $\tilde{\mathbb{E}}_{\infty}$ with exponential weights, i. e., $(z,\pi)=(v,Q,\ell,\omega,\pi)\in\tilde{\mathcal{K}}$ if and only if
(8.3) $\mathrm{e}^{\eta(\cdot)}(v,Q,\ell,\omega,\pi)\in\tilde{\mathbb{E}}_{\infty},\kern 5.0pt\text{with}\kern 5.0pt\|(v,Q,\ell,\omega,\pi)\|_{\tilde{\mathcal{K}}}\coloneqq\bigl\|\bigl(\mathrm{e}^{\eta(\cdot)}v,\mathrm{e}^{\eta(\cdot)}Q,\mathrm{e}^{\eta(\cdot)}\ell,\mathrm{e}^{\eta(\cdot)}\omega,\mathrm{e}^{\eta(\cdot)}\pi\bigr)\bigr\|_{\tilde{\mathbb{E}}_{\infty}}$.
For $\varepsilon>0$, the ball with center zero and radius $\varepsilon$ in $\tilde{\mathcal{K}}$ is denoted by $\tilde{\mathcal{K}}_{\varepsilon}$, and we define $\Psi\colon\tilde{\mathcal{K}}_{\varepsilon}\to\tilde{\mathcal{K}}$ to be the solution map to (8.1). This means that given $(z,\pi)=(v,Q,\ell,\omega,\pi)\in\tilde{\mathcal{K}}_{\varepsilon}$, we denote by $\Psi(z,\pi)$ the solution to (8.1) with $G_{1}(z,\pi)$, $G_{2}(z)$, $G_{3}(z,\pi)$ and $G_{4}(z,\pi)$. By Theorem 6.1(b), the map $\Psi$ is well-defined if the terms on the right-hand side lie in $\mathrm{e}^{-\eta(\cdot)}\mathbb{F}_{\infty}\coloneqq\{z=(v,Q,\ell,\omega)\in\mathbb{F}_{\infty}:\mathrm{e}^{\eta(\cdot)}z\in\mathbb{F}_{\infty}\}$, where $\mathbb{F}_{\infty}$ is defined in (6.1), and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, with $\mathrm{X}_{\gamma}$ from (2.10). Throughout the section, we use $\|\cdot\|_{\infty}$ to denote the $\mathrm{L}^{\infty}(0,\infty)$-norm, while $\|\cdot\|_{\infty,\infty}$ represents the norm of $\mathrm{L}^{\infty}(0,\infty;\mathrm{L}^{\infty}(\mathcal{F}_{0}))$.
In the lemmas below, we collect important ingredients for the nonlinear estimates. The first one discusses estimates of terms related to the transform. Section 8(a)–(c) and (e) follow from [9, Lemma 7.3 and Proposition 7.4]. Concerning Section 8(d), recall from [9, (7.17)] the estimate $\|\partial^{\beta}b_{i}\|_{\infty,\infty}\leq C\mathrm{e}^{-\eta t}\varepsilon$. In view of (5.2), this yields $\|\mathrm{J}_{X}-\operatorname{Id}_{3}\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}\leq C\varepsilon$ as well as $\|\mathrm{J}_{X}\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}\leq C$. The estimates of the second derivatives, of the terms involving $Y$ and of the other terms follow in a similar manner with regard to (5.3).
Lemma 8.1.
Recall $X$, $Y$ as well as $g^{ij}$, $g_{ij}$ and $\Gamma_{jk}^{i}$ from Section 5, and let $(v,Q,\ell,\omega,\pi)\in\tilde{\mathcal{K}}_{\varepsilon}$. Then
  1. (a)
    there exists an $\varepsilon$-independent constant $C>0$ with $\|\mathrm{J}_{X}-\operatorname{Id}_{3}\|_{\infty,\infty}\leq C\varepsilon$, so if $\varepsilon<\nicefrac{{1}}{{2C}}$, then $\mathrm{J}_{X}$ is invertible on $(0,\infty)\times\mathcal{F}_{0}$,
  2. (b)
    there is $C>0$ with $\|\mathrm{J}_{X}\|_{\infty,\infty}$, $\|\mathrm{J}_{Y}\|_{\infty,\infty}$, $\|g^{ij}\|_{\infty,\infty}$, $\|g_{ij}\|_{\infty,\infty}\leq C$ for all $i,j,k\in\{1,2,3\}$,
  3. (c)
    there exists $C>0$ so that for all $i,j,k\in\{1,2,3\}$, we have
    $\|\partial_{t}Y\|_{\infty,\infty}+\|\partial_{j}\partial_{t}X_{i}\|_{\infty,\infty}+\|\mathrm{J}_{Y}-\operatorname{Id}_{3}\|_{\infty,\infty}+\|\partial_{i}g^{jk}\|_{\infty,\infty}+\|\partial_{i}g_{jk}\|_{\infty,\infty}$
    $\quad+\|\Gamma_{jk}^{i}\|_{\infty,\infty}+\|g^{ij}-\delta_{ij}\|_{\infty,\infty}+\|\partial_{i}\partial_{j}X_{k}\|_{\infty,\infty}+\|\partial_{i}\partial_{j}Y_{k}\|_{\infty,\infty}\leq C\varepsilon,$
  4. (d)
    the assertions of (a)–(c) remain valid when $\|\cdot\|_{\infty,\infty}$ is replaced by $\|\cdot\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))}$, and
  5. (e)
    for $\mathbb{O}$ and $\mathbb{O}^{\top}$, we get $\|\mathbb{O}\|_{\infty}$, $\|\mathbb{O}^{\top}\|_{\infty}\leq C$ and $\|\mathbb{O}-\operatorname{Id}_{3}\|_{\infty}$, $\|\mathbb{O}^{\top}-\operatorname{Id}_{3}\|_{\infty}\leq C\varepsilon$.
In the following, we take into account small enough $\varepsilon_{0}$ such that $\mathrm{J}_{X}$ is invertible for all $\varepsilon\in(0,\varepsilon_{0})$ by Section 8(a). The next lemma provides estimates of some of the terms appearing in (8.2). Here Section 8(a) follows as in [9, Proposition 7.4], while for the estimates in (b), we invoke Section 8(d) and proceed as in the proof of Section 7 for the estimate of $(v\cdot\nabla)Q$. Concerning Section 8(c), we make use of $(\mathrm{T}-\mathcal{T})(v,\pi,Q)=(\nabla v+(\nabla v)^{\top})(\operatorname{Id}_{3}-\mathbb{O})-\frac{2\xi}{3}(\operatorname{Id}_{3}-\mathbb{O}^{\top})(\Delta Q-Q)-\frac{2\xi}{3}\mathbb{O}^{\top}(\Delta Q-Q)(\operatorname{Id}_{3}-\mathbb{O})$, recall the mapping properties of $\mathcal{J}$ from Section 7(b) and employ Section 8(e).
Lemma 8.2.
Let $p>4$, and for $\varepsilon\in(0,\varepsilon_{0})$, consider $(v,Q,\ell,\omega,\pi)\in\tilde{\mathcal{K}}_{\varepsilon}$. Then there exists $C>0$ so that
  1. (a)
    $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{L}_{1}-\Delta)v\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}\mathcal{M}v\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}\mathcal{N}(v)\|_{\mathbb{F}_{\infty}^{v}}$, $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{G}-\nabla)\pi\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon^{2}$,
  2. (b)
    $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{L}_{2}-\Delta)Q\|_{\mathbb{F}_{\infty}^{Q}}$, $\|\mathrm{e}^{\eta(\cdot)}(\partial_{t}Y\cdot\nabla)Q\|_{\mathbb{F}_{\infty}^{Q}}$, $\|\mathrm{e}^{\eta(\cdot)}(v\cdot\nabla)Q\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon^{2}$, and
  3. (c)
    for $\mathcal{J}$ as defined in Section 7(b), it is valid that $\|\mathrm{e}^{\eta(\cdot)}\mathcal{J}((\mathrm{T}-\mathcal{T})(v,\pi,Q))\|_{p}\leq C\varepsilon^{2}$.
We now successively prove the estimates of the nonlinear terms from (8.2).
Proposition 8.3.
Let $p>4$. Then there is $C>0$ so that for $\varepsilon\in(0,\varepsilon_{0})$ and $(z,\pi)$, $(z_{1},\pi_{1})$, $(z_{2},\pi_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, it holds that $\|\mathrm{e}^{\eta(\cdot)}G_{1}(z,\pi)\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon^{2}$ and $\|\mathrm{e}^{\eta(\cdot)}(G_{1}(z_{1},\pi_{1})-G_{1}(z_{2},\pi_{2}))\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon\cdot\|(z_{1},\pi_{1})-(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}$.
Proof.
By Section 8(a), it remains to estimate $\mathcal{B}_{\tau_{\mathrm{l}}}(Q)+\mathcal{B}_{\tau_{\mathrm{h}}}(Q)+\mathcal{B}_{\sigma}(Q)-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q$. We begin with $\mathcal{B}_{\sigma}(Q)$ from (5.7). The first two addends have the same structure as the others, so we concentrate on them. For this, we recall $\mathcal{L}_{2}$ from (5.4) and deduce from Section 8(c) that $\|\mathrm{e}^{\eta(\cdot)}\mathcal{L}_{2}Q\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon$. Arguing that the factor $Q_{ik}$ can be bounded by $C\varepsilon$ in $\mathrm{L}^{\infty}(0,\infty;\mathrm{L}^{\infty}(\mathcal{F}_{0}))$, estimating the first, second and third derivatives of $Q$ in $\mathrm{L}^{p}(0,\infty;\mathrm{L}^{2}(\mathcal{F}_{0}))$ with factor $\mathrm{e}^{\eta(\cdot)}$ by $\varepsilon$, and using Section 8 to bound the other factors, we obtain $\|\mathrm{e}^{\eta(\cdot)}\mathcal{B}_{\sigma}(Q)\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon^{2}$. The terms $\mathcal{B}_{\tau_{\mathrm{h},2}}$, $\mathcal{B}_{\tau_{\mathrm{h},3}}$, $\mathcal{B}_{\tau_{\mathrm{l},1}}$, $\mathcal{B}_{\tau_{\mathrm{l},2}}$ and $\mathcal{B}_{\tau_{\mathrm{l},3}}$ allow for analogous estimates thanks to the presence of at least two factors with estimates by $C\varepsilon$ in each addend.
For the estimate of the linear term $\mathcal{B}_{\tau_{\mathrm{h},1}}$ from (5.8), we invoke the term $\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q$. Recalling that it is given by $\nicefrac{{2\xi}}{{3}}\operatorname{div}(\Delta Q-Q)$, we expand the difference $\mathcal{B}_{\tau_{\mathrm{h},1}}(Q)-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q$ as
$-\frac{2}{3}\sum_{j,k=1}^{3}\Biggl(\sum_{l=1}^{3}\biggl(\sum_{m=1}^{3}\Bigl((\partial_{m}\partial_{l}\partial_{k}Q_{ij})(\partial_{j}Y_{m}-\delta_{jm})g^{kl}+(\partial_{m}\partial_{l}\partial_{k}Q_{ij})\delta_{jm}(g^{kl}-\delta_{kl})\Bigr)$
$\quad+(\partial_{l}\partial_{k}Q_{ij})(\partial_{j}g^{kl})+(\partial_{l}\partial_{k}Q_{ij})(\partial_{j}Y_{l})\Delta Y_{k}\biggr)+(\partial_{k}Q_{ij})(\partial_{j}\Delta Y_{k})+(\partial_{j}Y_{k}-\delta_{jk})(\partial_{k}Q_{ij})\Biggr).$
For estimates of $\partial_{j}Y_{m}-\delta_{jm}$ or $g^{kl}-\delta_{kl}$ by $C\varepsilon$, we exploit Section 8(c), while $Q$-terms yield an $\varepsilon$-power, and $\partial_{j}\Delta Y_{k}$ can be estimated by $C\varepsilon$ thanks to Section 8. The other terms can be bounded by Section 8, so $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{B}_{\tau_{\mathrm{h},1}}(Q)-\operatorname{div}S_{\xi}(0)(\Delta-\operatorname{Id}_{3})Q)\|_{\mathbb{F}_{\infty}^{v}}\leq C\varepsilon^{2}$. The Lipschitz estimates can be obtained in the same way upon modifying the above lemmas accordingly. ∎
Proposition 8.4.
Assume $p>4$. There exists $C>0$ so that for $\varepsilon\in(0,\varepsilon_{0})$, $(z,\pi)$, $(z_{1},\pi_{1})$, $(z_{2},\pi_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, we have $\|\mathrm{e}^{\eta(\cdot)}G_{2}(z)\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon^{2}$ and $\|\mathrm{e}^{\eta(\cdot)}(G_{2}(z_{1})-G_{2}(z_{2}))\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon\cdot\|(z_{1},\pi_{1})-(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}$.
Proof.
Thanks to Section 8(b), we only need to handle $\mathcal{S}(v,Q)+\tilde{S}_{\xi}(0)\nabla v$, $Q^{2}-\operatorname{tr}(Q^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}}$ and $\operatorname{tr}(Q^{2})Q$. The quadratic shape of the term $Q^{2}-\operatorname{tr}(Q^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}}$ implies that it can be estimated by $C\varepsilon^{2}$ upon invoking the first inequality in (7.7) which remains valid for $T=\infty$. The term $\operatorname{tr}(Q^{2})Q$ can be treated likewise.
For the estimates of the remaining terms, we first handle the term $\mathcal{D}_{ij}$ from (5.9). Indeed, making use of Hölder’s inequality and Section 8(d), we infer that $\|\mathrm{e}^{\eta(\cdot)}\mathcal{D}_{ij}\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon$. In view of the shape of $\mathcal{S}(v,Q)$ from (5.10) and estimates of the $Q$-factors by $\varepsilon$ in $\mathrm{L}^{\infty}(0,\infty;\mathrm{W}^{1,\infty}(\mathcal{F}_{0}))$ by virtue of Section 7(b), all terms in $\mathcal{S}(v,Q)$ except for $\nicefrac{{\xi}}{{3}}(\mathcal{D}_{ij}+\mathcal{D}_{ji})$ can thus be estimated by $C\varepsilon^{2}$. In order to guarantee suitable decay of this term, we invoke $\tilde{S}_{\xi}(0)\nabla v=-\nicefrac{{\xi}}{{3}}(\nabla v+(\nabla v)^{\top})$. We recall the representation of $\mathcal{D}_{v}=\mathcal{D}-\nabla v$ from (7.5), exploit Section 8(d) and observe that $\|\mathrm{e}^{\eta(\cdot)}\partial_{l}v_{k}\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{1}(\mathcal{F}_{0}))}\leq C\varepsilon$ to conclude $\|\mathrm{e}^{\eta(\cdot)}\mathcal{D}_{v}\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon^{2}$, and likewise for $\mathcal{D}_{v}^{\top}$. In total, we have $\|\mathrm{e}^{\eta(\cdot)}(\mathcal{S}(v,Q)+\tilde{S}_{\xi}(0)\nabla v)\|_{\mathbb{F}_{\infty}^{Q}}\leq C\varepsilon^{2}$. Note that the proof of the Lipschitz estimate is conceptually similar. ∎
Proposition 8.5.
Let $p>4$. Then there is $C>0$ so that for all $\varepsilon\in(0,\varepsilon_{0})$, $(z,\pi)$, $(z_{1},\pi_{1})$, $(z_{2},\pi_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$, it is valid that
$\|\mathrm{e}^{\eta(\cdot)}G_{3}(z,\pi)\|_{\mathbb{F}_{\infty}^{\ell}}$ $\leq C\varepsilon^{2}\kern 5.0pt\text{and}\kern 5.0pt\|\mathrm{e}^{\eta(\cdot)}(G_{3}(z_{1},\pi_{1})-G_{3}(z_{2},\pi_{2}))\|_{\mathbb{F}_{\infty}^{\ell}}\leq C\varepsilon\cdot\|(z_{1},\pi_{1})-(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}\kern 5.0pt\text{as well as}\kern 5.0pt$
$\|\mathrm{e}^{\eta(\cdot)}G_{4}(z,\pi)\|_{\mathbb{F}_{\infty}^{\omega}}$ $\leq C\varepsilon^{2}\kern 5.0pt\text{and}\kern 5.0pt\|\mathrm{e}^{\eta(\cdot)}(G_{4}(z_{1},\pi_{1})-G_{4}(z_{2},\pi_{2}))\|_{\mathbb{F}_{\infty}^{\omega}}\leq C\varepsilon\cdot\|(z_{1},\pi_{1})-(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}$.
Proof.
It readily follows that the terms $m_{\mathcal{S}}\omega\times\ell$ as well as $\omega\times(J_{0}\omega)$ can be estimated by $C\varepsilon^{2}$. With regard to (8.2) and Section 8, it remains to discuss the term $\int_{\partial\mathcal{S}_{0}}\mathbb{O}(t)^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O}(t)n\,\mathrm{d}\Gamma$ in $G_{3}$ and the analogous term in $G_{4}$. As in the proof of Section 7, we focus on selected terms. The other terms can be treated in the same way. We have already discussed the linear part of $\tau$ in Section 8, and we observe that every remaining term has at least two factors of terms with $Q$, allowing for an estimate by $C\varepsilon^{2}$. Similarly as in the proof of Section 7, with the respective embeddings applied with $T=\infty$, and employing Section 8(e) for the boundedness of $\mathbb{O}$ and $\mathbb{O}^{\top}$, for small $\delta>0$, we get $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}\nabla Q\odot\nabla Q\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\cdot\|Q\|_{\mathrm{L}^{\infty}(0,\infty;\mathrm{H}^{\nicefrac{{3}}{{2}}+\delta}(\mathcal{F}_{0}))}\cdot\|\mathrm{e}^{\eta(\cdot)}Q\|_{\mathrm{L}^{p}(0,\infty;\mathrm{H}^{\nicefrac{{3}}{{2}}+\delta}(\mathcal{F}_{0}))}\leq C\varepsilon^{2}$, $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}Q\operatorname{tr}(Q\Delta Q)\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\varepsilon^{3}$ and $\left\|\mathrm{e}^{\eta(\cdot)}\int_{\partial\mathcal{S}_{0}}(\mathbb{O}^{\top}Q\operatorname{tr}(Q^{2})\operatorname{tr}(Q^{2})\mathbb{O})n\,\mathrm{d}\Gamma\right\|_{p}\leq C\varepsilon^{5}$. Hence, recalling $\mathcal{J}$ from Section 7(b), we deduce that $\|\mathrm{e}^{\eta(\cdot)}\mathcal{J}(-\mathbb{O}^{\top}(\tau_{\mathrm{r}}(Q)+\sigma(Q))\mathbb{O})\|_{p}\leq C\varepsilon^{2}$, so the first part of the assertion is implied. The Lipschitz estimate can be obtained likewise. ∎
These preparations pave the way for the proof of the second main result.
Proof of Theorem 4.2.
First, we verify Theorem 5.2. We have already argued that $\Psi\colon\tilde{\mathcal{K}}_{\varepsilon}\to\tilde{\mathcal{K}}$ is well-defined. For $\varepsilon\in(0,\varepsilon_{0})$, with $\varepsilon_{0}>0$ as above, let $(z,\pi)$, $(z_{1},\pi_{1})$, $(z_{2},\pi_{2})\in\tilde{\mathcal{K}}_{\varepsilon}$. By $\|(v_{0},Q_{0},\ell_{0},\omega_{0})\|_{\mathrm{X}_{\gamma}}\leq\delta$, and exploiting Theorem 6.1(b) as well as Section 8, Section 8 and Section 8, we find
$\|\Psi(z,\pi)\|_{\tilde{\mathcal{K}}}\leq C_{\mathrm{MR}}(\delta+C\varepsilon^{2})\kern 5.0pt\text{and}\kern 5.0pt\|\Psi(z_{1},\pi_{1})-\Psi(z_{2},\pi_{2})\|_{\tilde{\mathcal{K}}}\leq C_{\mathrm{MR}}C\varepsilon.$
With $\delta\leq\nicefrac{{1}}{{2C_{\mathrm{MR}}}}$ and $\varepsilon\leq\min\{\varepsilon_{0},\nicefrac{{1}}{{2C_{\mathrm{MR}}C}},\nicefrac{{1}}{{2C}},\sqrt{\nicefrac{{\delta}}{{C}}}\}$, we conclude the self-map and contraction property of $\Psi$. Thus, there exists a unique fixed point $(z_{*},\pi_{*})=(v_{*},Q_{*},\ell_{*},\omega_{*},\pi_{*})$, being the unique global strong solution to (5.11)–(5.13). Theorem 5.2 then is a consequence of the definition of $\tilde{\mathcal{K}}$ in (8.3).
The last task is to deduce Theorem 4.2 from Theorem 5.2. The latter implies $(\ell,\omega)\in\mathrm{W}^{1,p}(0,\infty)^{6}$, so $h^{\prime}$, $\Omega$ and $X$ can be obtained as indicated in Section 5. For $v$, $Q$ and $\pi$, we invoke the backward change of coordinates from Section 5. By construction, this is the solution to (2.5)–(2.9) in the asserted regularity class. Finally, we prove that $\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}$ for all $t\in[0,\infty)$. The shape of $\mathcal{S}(t)$ at time $t$ as introduced in Section 2 and the above estimates imply that for all $t\in[0,\infty)$, we have
$\mathrm{dist}(\mathcal{S}(t),\mathcal{S}_{0})\leq|h(t)|+|\mathbb{O}(t)-\operatorname{Id}_{3}|<\nicefrac{{r}}{{2}},\kern 5.0pt\text{so}\kern 5.0pt\mathrm{dist}(\mathcal{S}(t),\partial\mathcal{O})>\nicefrac{{r}}{{2}}\kern 5.0pt\text{by}\kern 5.0pt\mathrm{dist}(\mathcal{S}_{0},\mathcal{O})>r.\qed$
Acknowledgements This research is supported by the Basque Government through the BERC 2022-2025 program and by the Spanish State Research Agency through BCAM Severo Ochoa excellence accreditation CEX2021-01142-S funded by MICIU/AEI/10.13039/501100011033 and through Grant PID2023-146764NB-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union. Felix Brandt would like to express his gratitude to the German National Academy of Sciences Leopoldina for support through the Leopoldina Fellowship Program with grant number LPDS 2024-07. Felix Brandt and Matthias Hieber acknowledge the support by DFG project FOR 5528. Arnab Roy would like to thank the Alexander von Humboldt-Foundation, Grant RYC2022-036183-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+. The authors would also like to thank the anonymous referees for the careful reading and valuable comments that helped improve the quality of the manuscript.